### v1

# Scalable LinUCB: Low-Rank Design Matrix Updates for Recommenders with Large Action Spaces

Abstract
Linear contextual bandits, especially LinUCB, are widely used in recommender systems. However, its training, inference, and memory costs grow with feature dimensionality and the size of the action space. The key bottleneck becomes the need to update, invert and store a design matrix that absorbs contextual information from interaction history. In this paper, we introduce Scalable LinUCB, the algorithm that enables fast and memory efficient operations with the inverse regularized design matrix. We achieve this through a dynamical low-rank parametrization of its inverse Cholesky-style factors. We derive numerically stable rank‑1 and batched updates that maintain the inverse without directly forming the entire matrix. To control memory growth, we employ a projector‑splitting integrator for dynamical low‑rank approximation, yielding average per‑step update cost $\mathcal{O}(dr)$ and memory $\mathcal{O}(dr)$ for approximation rank $r$. Inference complexity of the suggested algorithm is $\mathcal{O}(dr)$ per action evaluation. Experiments on recommender system datasets demonstrate the effectiveness of our algorithm.
Keywords: LinUCB algorithm, Low-rank approximation, Projector-Splitting integrator
1 Introduction
Interactive recommender systems must learn from sparse and delayed feedback while responding within tight latency budgets. A popular class of methods for solving this problem are the contextual bandits, which are gaining increasing popularity due to their ability to fine-tune on new data and efficiently manage the exploration–exploitation trade-off auer2002finite . These algorithms are not only well-studied theoretically, with strong finite-time guarantees auer2002finite , abbasi2011improved , agrawal2013thompson , but are also widely used in production recommender systems li2010contextual , yi2023online .
Within this class of methods, one of the most popular algorithms is the well-known LinUCB li2010contextual . The model is simple, its parameters are interpretable, and the update rule is easy to implement. LinUCB assumes a linear relationship between the $d$-dimensional context vector and the expected reward of the selected action. Regret guarantees for the LinUCB algorithm are well studied abbasi2011improved ; moreover, LinUCB admits practical implementations for online recommendations li2010contextual . These properties make it a reasonable baseline when a system requires online updates.
Scalability becomes an issue as the feature dimension $d$ and the number of items (actions) grow. Exact implementations Angioli2025Efficient maintain per-arm design matrices with ridge regularization and apply rank-1 updates using the Sherman–Morrison identity. This keeps training at $\mathcal{O}(d^{2})$ per round but still requires $\mathcal{O}(d^{2})$ memory per arm and inversion of the design matrix for each candidate to evaluate confidence widths (exploration bonuses). With large action sets, inference becomes prohibitively costly since all items (actions) must be scored per request. Prior works reduce these costs by sketching the design matrix (e.g., Frequent Directions and its compensated/adaptive variants) kuzborskij2019efficient , chen2020efficient , wen2024matrix , by projecting features to a lower dimension yu2017cbrap , or by using diagonal/block-diagonal approximations yi2023online . Each option involves a different trade-off between accuracy, robustness, and computational complexity.
In this paper, we propose a scalable version of the LinUCB algorithm based on dynamical updates of Cholesky-style symmetric factors of the inverse regularized design matrix. Our algorithm enables fast training and inference while preserving the quality of the original LinUCB algorithm. At the same time, it reduces the memory consumption and iteration cost of the method. Our primary contributions are as follows:
  • We maintain and dynamically update a Cholesky-style factor representation for the inverse regularized design matrix in the LinUCB algorithm. We derive the corresponding dynamical low-rank update, which avoids explicit computation of inverse matrices. We employ a projector-splitting method lubich2014projector for dynamical low-rank approximation to cap the approximation rank and obtain predictable time and memory costs, referring to the proposed method as PSI-LinUCB. The proposed algorithm yields an average per-interaction update cost of $\mathcal{O}(dr)$ during training, where $r$ is the approximation rank. The inference complexity of the PSI-LinUCB algorithm is $\mathcal{O}(dr)$ per action evaluation.
  • We evaluate our method on the classical MovieLens-1M dataset as well as on a set of large-scale recommendation datasets from the Amazon collection — the setting in which standard LinUCB immediately fails even on top-grade hardware due to scalability issues. The proposed method achieves a $8.8\times$ reduction in training and inference time with respect to the standard LinUCB algorithm with Sherman–Morrison updates Angioli2025Efficient . It also reduces memory consumption in $15.9\times$ while maintaining the quality metric (hit rate) at the level of the vanilla LinUCB implementation.
2 Problem setting
We begin this section with describing the LinUCB algorithm li2010contextual , arguably the most popular algorithm among the class of contextual bandits. This algorithm employs the Upper Confidence Bound (UCB) exploration strategy, following the principle of optimism in face of uncertainty auer2002finite . Theoretical properties of LinUCB algorithm attracted a vast amount of contributions abbasi2011improved , chu2011contextual , valko2013finite , lattimore2020bandit .
Formally, at each time step $t\in\{0,\ldots,T\}$, agent observes a set of arms (actions) $\mathcal{A}_{t}$, where each arm $a\in\mathcal{A}$ is associated with a context vector ${x}_{t,a}\in\mathbb{R}^{d}$. LinUCB assumes a linear relationship between the context vector and the reward, that is, expected reward for arm $a$ at time $t$ is given by
$\mathbb{E}[r_{t,a}|{x}_{t,a}]={x}_{t,a}^{\top}{\theta}_{a}^{*},$ (2)
where ${\theta}_{a}^{*}\in\mathbb{R}^{d}$ is the unknown true parameter vector for arm $a$. For each arm $a\in\mathcal{A}$, the algorithm maintains a ridge regression estimator with a design matrix
$A_{a}=\lambda I+\sum_{s=1}^{t-1}{x}_{s,a}{x}_{s,a}^{\top},$ (3)
reward vector
$b_{a}=\sum_{s=1}^{t-1}r_{s,a}{x}_{s,a},$ (4)
and parameter estimate
$\hat{{\theta}}_{a}=A_{a}^{-1}b_{a}.$ (5)
At time $t$, the algorithm selects the arm according to
$a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\hat{{\theta}}_{a}^{\top}{x}_{t,a}+\alpha\sqrt{{x}_{t,a}^{\top}A_{a}^{-1}{x}_{t,a}}\right),$ (6)
where $\alpha\in\mathbb{R}$ is a confidence parameter controlling the exploration–exploitation trade-off. In online learning problems $\alpha>0$, but sometimes learning problems with fixed dataset require setting negative values of $\alpha$ (the so-called anti-exploration), see e.g. rezaeifar2022offline , xu2024provably .
In the context of recommender systems, at each time step $t$, the agent observes a particular user $u$ and a set of available items $\mathcal{A}_{t}$. For consistency with the previous notation, the context vector $\mathbf{x}_{t,a}$ can be equivalently written as $\mathbf{x}_{u,a}$, representing the joint feature vector of the user–item pair $(u,a)$, which captures both user-specific and item-specific characteristics.
The LinUCB algorithm described above can be generalized to the setting with batch updates, where model parameters are updated after multiple interactions (after a batch of interactions) rather than once over the whole training dataset, see Algorithm˜1.
Algorithm 1 LinUCB
0: train_data = $batch_{0},\dots,batch_{n-1}]$, regularization parameter $\lambda$
1: for each arm $a\in\mathcal{A}$ set $A_{0,a}=\lambda I$
2:for t in {0,…,n-1} do
3:  $X_{t,a}=[]$, $R_{t,a}=[]$
4:  for all $(u,a,r)$ in $batch_{t}$ do
5:   $X_{t,a}.append\,[x_{u,a}]$
6:   $R_{t,a}.append\,[r]$
7:  end for
8:  for each arm $a$ in $batch_{t}$ do
9:   $A_{t+1,a}\leftarrow A_{t,a}+X_{t,a}^{\top}X_{t,a}$
10:   $b_{t+1,a}\leftarrow b_{t,a}+X_{t,a}^{\top}R_{t,a}$
11:   ${\theta}_{t+1,a}\leftarrow A_{t+1,a}^{-1}b_{t+1,a}$
12:  end for
13:end for
14:for each arm $a\in\mathcal{A}$ do
15:  Observe context $x_{u,a}$
16:  Compute $\text{UCB}_{u,a}={\theta}_{n,a}^{\top}{x}_{u,a}+\alpha\sqrt{{x}_{u,a}^{\top}{A}_{n,a}^{-1}{x}_{u,a}}$
17:end for
18:return $\arg\max_{a\in\mathcal{A}}\text{UCB}_{u,a}$
However, the LinUCB algorithm has several scalability challenges:
  • Matrix inversion: The inversion of the matrix $A_{a}$ has a computational complexity of $O(d^{3})$, which makes the algorithm computationally expensive in high-dimensional context spaces.
  • Large action space: A separate matrix $A_{a}\in\mathbb{R}^{d\times d}$ must be stored and updated for each arm, which not only increases time and memory requirements but also the number of matrix operations as the action space grows.
3 Related works
To address scalability constraints, several extensions and variants of the LinUCB algorithm have been proposed in the literature.
A widely used approach to accelerate LinUCB is to apply rank-1 updates of the ridge-regularized design matrix $A_{a}$ via the Sherman–Morrison identity  Angioli2025Efficient , Ciucanu2022Implementing , Wang2022Dynamic , Yan2025CoCoB , Ozbay2024Comparative . These updates reduce the per-round update complexity from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d^{2})$ while maintaining exact parameter estimates, but still require storing a full $d\times d$ matrix for each arm.
Another line of research aims to improve the efficiency of LinUCB through matrix sketching techniques. Early work kuzborskij2019efficient demonstrated that LinUCB, can be efficiently implemented using the Frequent Directions (FD) sketching method, reducing the per-round update time from $\mathcal{O}(d^{2})$ to $\mathcal{O}(md)$, where $m$ is the sketch size and $d$ is the feature dimension. However, applying FD to contextual bandits has certain drawbacks. In particular, the FD sketching method violate the positive definite monotonicity design matrices $A_{a}$, which may affect stability and theoretical guarantees. To address this limitation, a Spectral Compensation Frequent Directions (SCFD) method and its adaptation for high-dimensional (CBSCFD) algorithms were proposed chen2020efficient , which preserves positive definiteness while maintaining the same $\mathcal{O}(md)$ computational and memory complexity. More recently, adaptive sketching techniques such as Dyadic Block Sketching wen2024matrix have been introduced to dynamically adjust the sketch size, ensuring a per-round update complexity of $\mathcal{O}(dl)$, where $l$ is the current sketch size. This adaptive strategy prevents excessive spectral loss and avoids linear regret when the spectrum of the design matrix decays slowly.
Another perspective on improving the efficiency of linear contextual bandits is through dimensionality reduction via random projection. For example, the Contextual Bandits via Random Projection (CBRAP) algorithm yu2017cbrap , address the challenges of high-dimensional contexts by mapping the original $d$-dimensional features to a lower $m$-dimensional subspace. This reduces the per-round update complexity from $\mathcal{O}(d^{2})$ to $\mathcal{O}(md+m^{3})$.
One more approach to tackle the limitations of the LinUCB algorithm is to approximate each design matrix $A_{a}$ along its diagonal as proposed in Diag-LinUCB yi2023online . This allows scalable online updates running as $\mathcal{O}(d)$ per round in terms of both time and memory. As the authors demonstrate, in the specific contexts $x_{ua}$ such updates are sufficient to prevent the loss of model quality during training.
A detailed discussion of the limitations of the approaches reviewed above is provided in Section˜6.1.
4 Scalable LinUCB with low-rank updates
The main challenges related to LinUCB updates are its scalability with the number of arms $\mathcal{A}$ and the complexity of updating the parameter $\hat{\theta}_{a}$ in (5). Indeed, after observing each batch of interactions, the learner has to update the parameters that correspond to every arm $a$ appeared in the current batch. This update requires computing $A_{a}^{-1}$, which has the complexity of $\mathcal{O}(d^{3})$ in the naive implementation. Moreover, computing the decision-making rule (6) requires matrix-vector multiplications with the same matrix $A_{a}^{-1}$.
A natural idea to reduce complexity of calculating $A_{a}^{-1}$ is to consider the setting of rank-1 updates (when the context vector arrives one at a time) together with the Sherman-Morrison update equation. This approach was implemented in context of recommender systems in Angioli2025Efficient . Further ways to reduce the model complexity is to use a low-rank approximation of the matrix
$\sum_{s=1}^{t-1}{x}_{s,a}{x}_{s,a}^{\top}\in\mathbb{R}^{d\times d}\;,$ (7)
which is a part of $A_{a}$ in (3). This approach is used in the frequent directions (FD) algorithm kuzborskij2019efficient and CBSTD algorithm chen2020efficient . In contrast to these works, we employ an alternative approach, which relies on the dynamic representation of the inverse matrix $A_{a}^{-1}$. Towards this aim, we use the Cholesky-style symmetric factorization:
$A_{a,t}=L_{a,t}L_{a,t}^{\top}$ (8)
and note that in this case
$A_{a,t}^{-1}=L_{a,t}^{-\top}L_{a,t}^{-1}\;.$ (9)
The representation (9) is a key formula for our further analysis. We acknowledge the following important statement - we do not compute the inverse factors $L_{a,t}^{-1}$ from the matrix $L_{a,t}$, but use a dynamical updates for $L_{a,t}^{-1}$ based on $L_{a,t-1}^{-1}$. Since the update mechanism is identical for each arm, we omit the index $a$ and write $A_{t}$ and $L_{t}$ (respectively, $L_{t}^{-1}$) instead of $A_{a,t}$ and $L_{a,t}$.
In the next parts of this section, we present our methodology in different setups. First, in Section˜4.1 we consider the case of rank-1 updates and write the dynamics of updates of the inverse root matrix $L_{t}^{-1}$. Then in Section˜4.2 we generalize our expressions for the batch update case. Both representations rely on a recursive definition in the form of matrix decomposition $U_{t}V_{t}^{\top}$. In Section˜5, we show that this decomposition can be dynamically updated in a low-rank format using the Projector-Splitting Integrator (PSI) method of Lubich and Oseledets lubich2014projector . Such updating mechanism restrains an uncontrollable growth of factor matrices $U_{t}$ and $V_{t}$ over time and enables “on-the-fly” adaptation to the stream of contextual data. This approach has been previously considered in the context of modifying the PureSVD model in recommender systems Olaleke2021Dynamic .
4.1 Rank-1 updates
We start by considering the case of context vectors arriving one-by-one requiring the parameter updates to be performed after every interaction. The update to design matrix $A_{t}$ at time $t+1$ can be written in the form of a rank-1 perturbation
$A_{t+1}=A_{t}^{\top}+x_{t+1}x_{t+1}^{\top}.$ (10)
A straightforward approach to calculating updates to the inverse is to apply the Sherman-Morrison formula and obtain the recursive form through $t$ as done by Angioli2025Efficient . We propose a different representation that enables more efficient computations via our low-rank update technique.
We first rewrite the rank-1 perturbation in the Cholesky-style matrix decomposition form as follows:
$A_{t+1}=L_{t}L_{t}^{\top}+x_{t+1}x_{t+1}^{\top}=L_{t}(I+L_{t}^{-1}x_{t+1}x_{t+1}^{\top}L_{t}^{-\top})L_{t}^{\top}.$ (11)
Then we can rewrite the design matrix as
$A_{t+1}=L_{t+1}L_{t+1}^{\top}\;,$ (12)
where we set
$L_{t+1}=L_{t}(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})\;,\quad\bar{x}_{t+1}=L_{t}^{-1}x_{t+1}\;,$ (13)
and parameter $\alpha_{t+1}\in\mathbb{R}$ such that
$1+\alpha_{t+1}\,\|\bar{x}_{t+1}\|^{2}\;=\;\sqrt{\,1+\|\bar{x}_{t+1}\|^{2}\,}\;.$ (14)
The correctness of the specified recursive definition of $L_{t+1}$ in (13) can be verified by direct substitution. The following proof for the rank-1 updates setting will be useful for illustrating the essence of our approach.
Theorem 1.
Let $\varepsilon>0$, $L_{0}=\varepsilon^{-1/2}I\in\mathbb{R}^{d\times d}$, and $U_{0}$, $V_{0}$ be empty matrices. Given a sequence of context vectors $\{x_{t}\}_{t\in\mathbb{N}}$ set
$\beta_{t+1}\;=\;\frac{\alpha_{t+1}}{1+\alpha_{t+1}\,\|\bar{x}_{t+1}\|^{2}}.$ (15)
Then the inverse root $L_{t+1}^{-1}$ can be expressed as
$L^{-1}_{t+1}\;=\;\bigl(I-U_{t+1}\,V_{t+1}^{\top}\bigr)\,L_{0}^{-1},$ (16)
where the matrices $U_{t+1}\in\mathbb{R}^{d\times(t+1)}$ and $V_{t+1}\in\mathbb{R}^{d\times(t+1)}$ are recursively updated with column-wise concatenation
$U_{t+1}\;=\;\begin{bmatrix}U_{t}&\beta_{t+1}\,\bar{x}_{t+1}\end{bmatrix},\qquad V_{t+1}\;=\;\begin{bmatrix}V_{t}&(I-V_{t}\,U_{t}^{\top})\,\bar{x}_{t+1}\end{bmatrix}$. (17)
Proof.
Using (13), we obtain
$L_{t+1}^{-1}=(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})^{-1}L_{t}^{-1}\;.$ (18)
After applying the Sherman-Morrison-Woodbury formula
$(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})^{-1}=I-\frac{\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}}{1+\alpha_{t+1}\bar{x}_{t+1}^{\top}\bar{x}_{t+1}}=I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}\;,$ (19)
the recursive structure of the inverse update reads:
$L_{t}^{-1}=(I-\beta_{t}\bar{x}_{t}\bar{x}_{t}^{\top})L_{t-1}^{-1}\;.$ (20)
Unraveling this recursion from $t$ back to the initial condition gives
$L_{t}^{-1}=(I-\beta_{t}\bar{x}_{t}\bar{x}_{t}^{\top})L_{t-1}^{-1}=\prod_{i=1}^{t}(I-\beta_{i}\bar{x}_{i}\bar{x}_{i}^{\top})L_{0}^{-1}.$ (21)
We now prove by induction that $L^{-1}_{t}=(I-U_{t}\,V_{t}^{\top})L^{-1}_{0}$. Indeed, for $t=1$, we have
$L_{1}^{-1}=(I-\beta_{1}\bar{x}_{1}\bar{x}_{1}^{\top})L_{0}^{-1}=(I-U_{1}V_{1}^{\top})L_{0}^{-1}\;,$ (22)
where $U_{1}=\beta_{1}\bar{x}_{1}$ and $V_{1}=\bar{x}_{1}$. Assuming that $L^{-1}_{t}=(I-U_{t}\,V_{t}^{\top})L^{-1}_{0}$,
$L_{t+1}^{-1}$ $=(I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})L_{t}^{-1}$ (23)
$=(I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})(I-U_{t}\,V_{t}^{\top})L^{-1}_{0}$ (24)
$=(I-U_{t}V_{t}^{\top}-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}+\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}U_{t}V_{t}^{\top})L_{0}^{-1}$ (25)
$=(I-(U_{t}V_{t}^{\top}+\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}(I-U_{t}V_{t}^{\top})))L_{0}^{-1}\;.$ (26)
This can be written as $L_{t+1}^{-1}=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1}$, where
$U_{t+1}$ $=\begin{bmatrix}U_{t}&\beta_{t+1}\,\bar{x}_{t+1}\end{bmatrix}$, (27)
$V_{t+1}$ $=\begin{bmatrix}V_{t}&(I-V_{t}U_{t}^{\top})\bar{x}_{t+1}\end{bmatrix}$. (28)
This completes the proof. ∎
4.2 Batch updates
Now we provide a generalization of Theorem˜1 to the case of batch updates. In this setting parameter is updated after some number of interactions is accumulated in the system. Formally, we write $X_{t}\in\mathbb{R}^{d\times B}$ for the batch of $B\in\mathbb{N}$ concatenated contexts collected during the $t$-th interaction round with the hand $a\in\mathcal{A}$. Note that, generally speaking, $B$ depends on $a$ and $t$, but we prefer to write $B$ instead of $B_{t,a}$ for notation simplicity.
The update rule for the regularized design matrix can then be expressed as
$A_{t+1}=A_{t}+X_{t+1}X_{t+1}^{\top}=L_{t+1}L_{t+1}^{\top},$ (29)
where the Cholesky-like factor $L_{t+1}$ is obtained recursively as
$L_{t+1}=L_{t}(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}),$ (30)
with $Q_{t+1}\in\mathbb{R}^{d\times B}$ being a matrix with orthonormal columns and $Y_{t+1}\in\mathbb{R}^{B\times B}$ obtained using the fast symmetric factorization approach Ambikasaran described in Appendix˜B (see Theorem˜3).
Theorem 2.
Let $\varepsilon>0$, $L_{0}=\varepsilon^{-1/2}I\in\mathbb{R}^{d\times d}$ and $U_{0},V_{0}$ be empty matrices, and $U_{t},V_{t}\in\mathbb{R}^{d\times d_{t}}$. Then the factor $L_{t+1}$ defined in (30) can be updated by formula
$L_{t+1}^{-1}=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1}.$ (31)
Matrices $U_{t+1},V_{t+1}\in\mathbb{R}^{d\times(d_{t}+B)}$ are updated with column-wise concatenation
[…]
Proof of Theorem˜2 is provided in Section˜B.2. The shape $d_{t}$ of the matrices $U_{t}$ and $V_{t}$ can be inferred from the representations (31) - (2).
Note that the matrices $U_{t}$ and $V_{t}$ in (2) expand in the number of columns with time $t$, leading to excessive memory demands and significant computational overhead. At the same time, in practice, the context distributions associated with each action typically have low-rank structure. Below we show, how one can maintain the low-rank structure of $U_{t}$ and $V_{t}$, using the Projector-Splitting Integrator lubich2014projector .
5 Projector-Splitting Integrator
Now we aim to improve the LinUCB algorithm based on the representation (31)-(2). As already mentioned, the matrices $U_{t}$ and $V_{t}$ here grow with time $t$, which yields both computational and memory issues. At the same time, the key element of the representation (31) is the time-varying factor $D_{t}=U_{t}V_{t}^{\top}\in\mathbb{R}^{d\times d}$. In this case, we essentially observe the matrix dynamics $\{D_{t}\}_{t\in\mathbb{N}}$. Moreover, the matrices $D_{t}$ might be well-approximated by low-rank factors. One of the options for dynamic approximation of $D_{t}$ comes from a singular value decomposition
$D_{t}=\bar{U}_{t}\Sigma_{t}\bar{V}_{t}^{\top}\;.$ (32)
Here the best rank-$r$ approximation is given by the truncated SVD. At the same time, the orthogonal factors $(\bar{U}_{t},\bar{V}_{t})$ and $(\bar{U}_{t+1},\bar{V}_{t+1})$, corresponding to the matrices $D_{t}$ and $D_{t+1}$, are not guaranteed to be close, which might yield additional computational instability. Instead, we propose in this paper to rely on the projector-splitting integrator (PSI) approach of lubich2014projector , which aims at a dynamical low-rank approximation objective
$\|\dot{D}_{t}-\dot{\tilde{D}}_{t}\|\to\min_{\tilde{D}_{t}:\,\operatorname{rank}{\tilde{D}}_{t}=r}\,$, (33)
where $\dot{D}_{t}=\frac{\mathrm{d}}{\mathrm{d}t}D_{t}$. The PSI approach enables efficient updates to the approximations $\tilde{D}_{t}$ in an iterative manner, without explicitly forming the matrices $\tilde{D}_{t}$. This algorithm maintains dynamical form of updates
$\tilde{D}_{t}=\tilde{U}_{t}S_{t}\tilde{V}_{t}^{\top}\;,$ (34)
where the factors $\tilde{U}_{t},\tilde{V}_{t}\in\mathbb{R}^{d\times r}$ have orthonormal columns, and $S_{t}\in\mathbb{R}^{r\times r}$ is invertible. The matrix $S_{t}$ here can be diagonal, yet it is not required. Then, with computational complexity of QR decompositions of $\mathbb{R}^{d\times r}$ matrices and multiplication of matrices of appropriate shapes, we obtain the factors $(\tilde{U}_{t+1},S_{t+1},\tilde{V}_{t+1}^{\top})$, such that
$\tilde{D}_{t+1}=\tilde{U}_{t+1}S_{t+1}\tilde{V}_{t+1}^{\top}\;.$ (35)
We provide the implementation details in Algorithm˜4.
Now we explain the main steps of integrating Algorithm˜4 into the LinUCB algorithm with batch updates. At each batch update step, two matrices, $U_{t}$ and $V_{t}$, are maintained to represent the low-rank structure. Initially, these matrices are incrementally expanded by (27) or (2) as new data arrive, until their number of columns exceed a predefined rank threshold $r$. Once this occurs at some time step $t_{0}$, we compute the singular value decomposition (SVD) with fixed rank $r$ of the current product $U_{t_{0}}V_{t_{0}}^{\top}=\tilde{U}_{t_{0}}S_{t_{0}}\tilde{V}_{t_{0}}^{\top}$ to obtain the new factors $U_{t_{0}}=\tilde{U}_{t_{0}}S_{t_{0}}$, and $V_{t_{0}}=\tilde{V}_{t_{0}}$ for initialization Algorithm˜4. In subsequent batch updates, the matrices $U_{t}$ and $V_{t}$ are efficiently updated using the Algorithm˜4, which maintains the r-rank approximation. The corresponding increments $\Delta D_{t+1}$ defined by the algorithm’s update mechanism $U_{t+1}V_{t+1}^{\top}=U_{t}V_{t}^{\top}+\Delta D_{t+1}$ are directly obtained from (10) for the rank-1 update case:
$\Delta D_{t+1}=\beta_{t+1}\tilde{x}_{t+1}\tilde{x}_{t+1}^{\top}(I-U_{t}V_{t}^{\top});$ (36)
and from (2) for the batch update case:
$\Delta D_{t+1}=Q_{t+1}(Y_{t+1}^{-1}+I)^{-1}Q_{t+1}^{\top}(I-U_{t}V_{t}^{\top}).$ (37)
The complete integration of this procedure within the LinUCB algorithm is summarized in Algorithm˜2. We highlight that we do not need to form the matrices $L_{t}^{-1}$ given in (31) explicitly.
Complexity analysis. Training stage. We analyze the computational complexity of the proposed PSI-LinUCB algorithm and we start from its training phase (described in Algorithm˜2). The computational complexity of PSI step
$U_{t+1},V_{t+1}=\textit{PSI}(U_{t},V_{t},\;\Delta D_{t+1})$ (38)
does not depend on the number of interactions $B=B_{t,a}$ (that is, interactions with arm $a$ inside batch number $t$) and scales as $\mathcal{O}(dr^{2})$. This factor comes from the QR decomposition applied to $\mathbb{R}^{d\times r}$ matrices. Note that the complexity of computing SVD (line $11$ in Algorithm˜3) is also $\mathcal{O}(d(r+B)^{2}+(r+B)^{3})$. This can be achieved by computing the QR decomposition of the factors $U_{t},V_{t}$ of size at most $\mathbb{R}^{d\times(r+B)}$. The second dimension up to $r+B$ is due to line 10 in Algorithm˜3. Moreover, this SVD needs to be computed only once per each arm $a\in\mathcal{A}$. The last step contributing to asymptotic complexity is line $5$ in Algorithm˜3 with the computational complexity $\mathcal{O}(dB^{2}+B^{3})$ due to complexity of the QR and Cholesky decompositions. Thus, the overall complexity of handling blocks in the Update_arm algorithm (Algorithm˜3) is
$\mathcal{O}(d(r^{2}+B^{2})+r^{3}+B^{3})\;,$ (39)
as soon as this arm has accumulated at least $r$ interactions.
Inference stage. To compute the bonus term in (6) we use the decomposition defined earlier:
$\sqrt{x_{u,a}^{\top}A_{t}^{-1}x_{u,a}}=\sqrt{x_{u,a}^{\top}(L_{t}L_{t}^{\top})^{-1}x_{u,a}}=\|(I-U_{t}V_{t}^{\top})x_{u,a}\|$ (40)
Then at time $t$, the arm is selected by formula:
$\begin{split}a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\theta_{a}^{\top}x_{u,a}+\alpha\|(I-U_{t}V_{t}^{\top})x_{u,a}\|\right)\;.\end{split}$
The computational complexity of this operation is $\mathcal{O}(dr)$ per arm.
Remark 1.
The primary computational cost of our algorithm arises during the training phase, which requires on average $\mathcal{O}(d(r^{2}/B+B)+r^{3}/B+B^{2})$ operations per interaction. By choosing $B\propto r$ and assuming $r\ll d$, this reduces to average complexity of $\mathcal{O}(dr)$ operations per iteration. To our knowledge, the $\mathcal{O}(dr)$ average complexity is the best computational cost achieved by LinUCB-based algorithms. This complexity is achieved by the algorithms suggested in the previous papers chen2020efficient , wen2024matrix .
Algorithm 2 PSI-LinUCB (LinUCB training with PSI)
0: train_data = $batch_{0},\dots,batch_{n-1}]$, batch_size, U_0,V_0U\_0,\,V\_0
[…]
2:for t in {0,…,n-1} do
3:  $X_{t,a}=[]$, $R_{t,a}=[]$
4:  for all $(u,a,r)$ in $batch_{t}$ do
5:   $X_{t,a}.append\,[x_{u,a}]$
6:   $R_{t,a}.append\,[r]$
7:  end for
8:  for each arm $a$ in $batch_{t}$ do
9:   $U_{t+1,a},V_{t+1,a},\theta_{t+1,a}\leftarrow$ Update_arm($a$)
10:  end for
11:end for
12:return $\theta_{n,a}$, $U_{n,a},V_{n,a}$ for each arm $a\in\mathcal{A}$.
Algorithm 3 Update_arm($a$)
[…]
1: Set $b_{t}=b_{t,a}$, $U_{t}=U_{t,a}$, $V_{t}=V_{t,a}$, $X_{t}=X_{t,a}$, $R_{t}=R_{t,a}$, {Omitting index $a$ for simplicity}
[…]
6:if $U_{t}.\text{shape}[1]<r$ then
7:  $U_{t+1}\;\leftarrow\;[\,U_{t},\;Q_{t+1}C_{t+1}]$
8:  $V_{t+1}\;\leftarrow\;[\,V_{t},\,(I-V_{t}U_{t}^{\top})\;Q_{t+1}\,]$
9:else
10:  if first time $U_{t}.\text{shape}[1]\geq r$ then
11:   $\tilde{U}_{t+1}S_{t+1}\tilde{V}_{t+1}^{\top}=SVD(U_{t}V_{t}^{T})$
12:   $U_{t+1}\;\leftarrow\;\tilde{U}_{t+1}S_{t+1}$
13:   $V_{t+1}\;\leftarrow\;\tilde{V}_{t+1}$
14:  else
15:   $\Delta D_{t+1}=\,Q_{t+1}\,C_{t+1}\,Q_{t+1}^{\top}\,(I-U_{t}V_{t}^{\top}),$ {We do not form $\Delta D_{t+1}$ explicitly}
16:   $U_{t+1},V_{t+1}=\textit{PSI}(U_{t},V_{t},\;\Delta D_{t+1})$
17:  end if
18:end if
[…]
21:return $U_{t+1},V_{t+1}$, $\theta_{t+1}$
Algorithm 4 PSI(Projector-Splitting Integrator)
[…]
Algorithm 5 Calculate_C_and_Q
[…]
4:Y=MIY=M-I// $Y\in\mathbb{R}^{B\times B}$
[…]
6:return $C_{t+1},Q_{t+1}$
6 Experimental Setup
Our experiments are conducted on four real-world datasets from Amazon product recommendations and classic Movielens 1M. In Sections 7.1 and 7.2, we demonstrate the efficiency of our algorithm compared to the batch LinUCB and LinUCB Classic for large context and arm spaces. In Section 7.3, we show that our algorithm can match the baseline performance without storing full-rank matrices, using only a small rank instead. In Section 7.4, we demonstrate its quality advantage over a CBSBFD algorithm with similar computational complexity.
Our results are conducted on $4$ real-world datasets from Amazon collection and the classical Movielens 1M dataset. Information about datasets is summarized in Table˜1.
All experiments are conducted on an Intel(R) Xeon(R) Gold 6348 CPU, 2.60GHz, kernel: 5.4.0-96-generic.
6.1 Baseline selection
We evaluate the performance of our proposed PSI-LinUCB algorithm against three baselines: the batch LinUCB, LinUCB with rank-1 updates via the Sherman–Morrison formula (LinUCB Classic), and the CBSCFD method proposed in chen2020efficient (see a detailed description of the CBSCFD method in Section˜3). Other approaches have several drawbacks, and we discuss below the reasons why we did not choose these algorithms as baselines.
The Frequent Directions (FD) sketching method kuzborskij2019efficient violates the positive-definite monotonicity property of the design matrices $A_{a}$, which may affect the algorithm’s stability. Therefore, we compare our approach against CBSCFD chen2020efficient , which addresses these limitations while maintaining the same $\mathcal{O}(md)$ computational and memory complexity. Similarly, chen2020efficient shows that CBSCFD outperforms both the CBRAP (Contextual Bandits via Random Projection) algorithm of yu2017cbrap and the Frequent Directions method of kuzborskij2019efficient . For this reason, we compare directly only with the stronger baseline method.
Dyadic Block Sketching wen2024matrix is an adaptive sketching technique whose theoretical advantages rely on restrictive assumptions about the spectrum of the design matrix, which may not hold in real-world data. Furthermore, in practice, adaptive sketching methods require choosing the initial sketch size $l_{0}$ and the error tolerance $\epsilon$ (see Algorithm in wen2024matrix ), and the performance of the algorithm is reported to be very sensitive to these hyperparameters. Specifically, under unfavorable spectral conditions, the sketch may grow excessively, erasing computational gains, whereas small values of $\epsilon$ degrade accuracy. This makes the approach operationally less stable than CBSCFD chen2020efficient . Moreover, the authors of wen2024matrix did not apply their method to real-world recommender system benchmarks, focusing only on classification problems.
Finally, Diag-LinUCB yi2023online is applicable only to contexts generated by a particular embedding model and cannot be applied to arbitrary feature representations. The method also reduces the context dimensionality to control latency, while our experiments focus on scalability with respect to high-dimensional contextual features.
Table 1: Information about the datasets used for validation
Dataset #Users #Items #Interactions Density
Magazine Subscriptions 60,100 3,400 71,500 0.35%
Health & Personal Care 461,700 60,300 494,100 0.02%
All Beauty 632,000 112,600 701,500 0.01%
Musical Instruments 1,800,000 213,600 3,000,000 0.001%
Movielens 1M 6,040 3,706 1,000,209 4.18%
6.2 Training Protocol
We divide the training process into phases:
  1. 1.
    Warm-up phase: Initial training on 80% of historical data
  2. 2.
    Online training phase: Sequential learning on the remaining 20% of data, divided into equal-sized temporal intervals simulating monthly updates. The model is incrementally trained on each month’s data.
  3. 3.
    Evaluation: Model evaluated after each monthly training update to assess recommendation quality over time.
6.3 Feature Extraction
User and item contexts are extracted using SVD to obtain. Given an interaction matrix $A\in\mathbb{R}^{m\times n}$ with $m$ users and $n$ items, we apply truncated SVD with rank $r^{\prime}=\min(r,\min(m,n)-1)$ to decompose it as $A\approx U\Sigma V^{T}$.
Item features are directly obtained from the right singular vectors after sorting by decreasing singular values: $V^{\top}\in\mathbb{R}^{n\times r}$. User features are then computed through matrix multiplication: $A\cdot V^{\top}\in\mathbb{R}^{m\times r}$.
SVD features are extracted from the data available during the warm-up period, which corresponds to the initial training split. To guarantee that each user and item in the test set has a valid context representation derived from this phase, all cold users are excluded from evaluation. This restriction arises because SVD-based feature extraction requires an existing interaction history. The use of SVD features enables the construction of high-dimensional contextual embeddings required for the experiments, however, our algorithm do not require the specific context representation and is capable of operating with arbitrary context.
6.4 Context Representation
The LinUCB algorithm is implemented with a batch processing approach for scalable recommendation systems. The algorithm maintains hybrid feature vectors ${x}_{ua}$ by taking the outer product of the user features $x_{u}\in\mathbb{R}^{d_{u}}$ and the item features ${x}_{a}\in\mathbb{R}^{d_{a}}$, resulting in $d=(r^{\prime})^{2}=\mathbb{R}^{d_{u}\times d_{a}}$ dimensional contexts.
7 Results
7.1 Scalability with context size
This experiment evaluates how the algorithms scale with the increasing context $d$ from 100 to 12,000.
We use the Amazon Magazine Subscriptions dataset, as it combines realistic user–item interactions with a moderate number of arms and measure computational time (training and evaluation) and memory usage while varying the SVD rank $r^{\prime}$ on a logarithmic scale. During the evaluation phase, we average the time and memory usage across months, as each month contains approximately the same number of interactions (see Section˜6.2), making it reasonable to consider the average time per month.
As shown in Figure˜1, the LinUCB requires $\mathcal{O}(d^{3})$ time for the warm-up phase, $\mathcal{O}(d^{2})$ time for inference, and $\mathcal{O}(d^{2})$ memory, since it maintains the full matrix $A_{a,t}$ for each arm $a$.
LinUCB Classic has both time and memory complexity of $\mathcal{O}(d^{2})$. The lower time complexity, compared to LinUCB, is achieved through rank-1 updates using the Sherman–Morrison–Woodbury formula.
In contrast, PSI achieves complexity $\mathcal{O}(dr^{2})$ time for the warm-up (see Remark˜1), $\mathcal{O}(dr)$ time for inference, and $\mathcal{O}(dr)$ memory. This reduction is possible because PSI does not store the full matrix $A_{a,t}$, but only its dynamically updated low-rank factors.
Figure 1: Performance comparison across different context sizes on Magazine Subscriptions dataset .
7.2 Scalability with the number of arms
This experiment evaluates how the algorithms scale with the increasing number of arms.
We fix the optimal context dimensions for each dataset (Magazine, Health, Beauty, and Musical Instruments) using outer-product features selected by scree plot (see Section˜A.1.1) and choose the optimal PSI rank by cross-validaton (see Section˜A.1.1)). The number of arms is gradually increased from a small subset to the full action space. At each stage, for each state we select items with the highest observation frequency. Hyperparameters are tuned via cross-validation for each dataset and model (see Section˜A.1.2).
We compare PSI with two baselines, LinUCB and LinUCB Classic, measuring hit rate, computation time, and memory usage. As shown in Table˜3 and Figure˜27, both baseline methods fail to run at larger scales due to memory constraints, while PSI remains stable and achieves comparable perfomance with significantly lower time and memory costs.
Figure 2: Algorithm scaling with number of arms on Beauty dataset.
7.3 Recommendation quality vs. rank
To analyze how the rank parameter affects recommendation quality, we conducted an experiment in which we varied the rank of the PSI method while keeping the context dimensions and number of arms fixed for each dataset. We measured the quality of recommendations after the warm-up phase. At each rank configuration, we measured the hit rate and compared it against the LinUCB baseline. Each month we continued to train the model further and then measured its performance by calculation the relative hit rate to provide a demonstration of the improvement in the quality of our algorithm.
The experiment was conducted using the optimal hyperparameters (see  Section˜A.1.1 and Section˜A.1.2).
The results show that, starting from a relatively low rank, our method performs well across all datasets. This indicates that our approach is capable of maintaining high-quality recommendations without requiring a full-rank matrix.
Refer to caption Refer to caption
(a) MovieLens 1M (b) Amazon All Beauty
Refer to caption Refer to caption
(c) Amazon Health (d) Magazine Subscriptions
Figure 3: Quality comparison by months: PSI vs LinUCB
7.4 CBSCFD vs PSI
This experiment compares the performance of the CBSCFD and PSI-LinUCB algorithms across multiple datasets. In the original CBSCFD method, a shared design matrix $A$ is used for all arms, which does not allow personalization. We extend this approach by training a separate matrix $A_{a}$ for each arm $a$, as proposed in li2010contextual .
CBSCFD supports only rank-1 updates, where a single context arrives at each time step, while PSI-LinUCB generalizes to support batch-context updates, where multiple contexts can be accumulated and updated at once.
Table 2 presents the hit rate performance across four datasets (MovieLens 1M, Amazon All Beauty, Amazon Health, and Magazine Subscriptions) for both algorithms with varying parameter configurations. PSI-LinUCB consistently outperforms CBSCFD.
Figure 4 shows the relationship between the parameter $m$ and hit rate performance for CBSCFD across the four real-world datasets. The results show significant variation in the optimal $m$ values across different datasets, with unpredictable behavior. Even small changes in $m$ lead to varying results.
Figure 5 provides a corresponding analysis for PSI-LinUCB, demonstrating a consistent improvement in performance as the rank increases. The algorithm shows predictable behavior: the hit rate increases monotonically with rank. This trend holds across all datasets, including sparse matrices. The relationship between rank and performance remains stable, independent of dataset characteristics.
In the CBSCFD algorithm, the parameter $\alpha$ (which corresponds to $\varepsilon$ in our notations) is dynamically updated at each round: the initial value is obtained by accumulating the sum of lost singular values that are subtracted during matrix compression. This works as adaptive regularization that compensates for the lost information and maintains the positive definiteness of the approximation matrices. In our PSI-LinUCB algorithm, the regularization parameter $\varepsilon$ is fixed to 1; making it dynamic could improve stability by keeping the design matrix better conditioned, as classical LinUCB is sensitive to this parameter (the algorithms have the same optimal hyperparametres).
Table 2: Hit rate performance across datasets and algorithms
Dataset M1 M2 M3 M4 M5 M6 M7 M8 Avg
MovieLens 1M
PSI-LinUCB 0.065 0.102 0.075 0.054 0.041 0.046 0.048 0.046 0.060
CBSCFD 0.065 0.102 0.075 0.054 0.041 0.045 0.045 0.040 0.058
Amazon All Beauty
PSI-LinUCB 0.012 0.009 0.022 0.015 0.014 0.003 0.011 0.008 0.012
CBSCFD 0.008 0.010 0.022 0.017 0.019 0.002 0.011 0.011 0.012
Amazon Health
PSI-LinUCB 0.044 0.024 0.018 0.018 0.037 0.01 0.017 0.012 0.022
CBSCFD 0.044 0.024 0.014 0.018 0.037 0.01 0.01 0.007 0.020
Magazine Subscriptions
PSI-LinUCB 0.603 0.528 0.578 0.504 0.570 0.605 0.565 0.286 0.530
CBSCFD 0.595 0.538 0.586 0.496 0.561 0.613 0.548 0.257 0.524
Refer to caption Refer to caption
(a) MovieLens 1M (b) Amazon All Beauty
Refer to caption Refer to caption
(c) Amazon Health (d) Magazine Subscriptions
Figure 4: Average Hit rate for different $m$ for CBSCFD.
Refer to caption Refer to caption
(a) MovieLens 1M (b) Amazon All Beauty
Refer to caption Refer to caption
(c) Amazon Health (d) Magazine Subscriptions
Figure 5: Average Hit rate for different rank values for PSI-LinUCB.
8 Conclusion
We presented Scalable-LinUCB, a computationally efficient variant of LinUCB that maintains a dynamically low-rank representation of the inverse Cholesky-style symmetric factor of the regularized design matrix. The method allows to control rank growth of the accumulated context matrix with a projector–splitting step. Moreover, our method has an average complexity of updates of order $\mathcal{O}(dr)$ per candidate. One of the potential future research directions involves substituting the projector–splitting integrator with other dynamical low-rank updates.
Acknowledgements
This research was supported in part through computational resources of HPC facilities at HSE University kostenetskiy2021hpc .
Appendix A Experiments
A.1 Hyperparameter selection
A.1.1 Context dimensionality $r^{\prime}$.
The optimal values of $r^{\prime}$ for context size were selected using the scree plot, which displays the singular values in descending order, indicating the optimal number of components that capture the most informative structure of the data.
Dataset Optimal $r^{\prime}$ Context size PSI rank $r$
Movielens 1M 13 169 32
Magazine Subscriptions 41 1681 64
Health & Personal Care 37 1369 128
All Beauty 51 2601 128
Musical Instruments 70 4900 256
A.1.2 Regularization $\alpha$.
We used a grid search to find the optimal exploration parameter $\alpha$. First, a broad grid
$\alpha\in\{-10,-8,-6,-4,-2,0,2,4,6,8,10\}\,\,(\text{step size}\,2)$ (41)
was evaluated to assess the overall impact of different values of $\alpha$ on the model’s performance. Based on the results from this initial search, a more refined range was selected:
$\alpha=\{-0.1,\,-1,\,-0.5,\,-0.3,\,-0.2,\,-0.05,\,-0.03,\,-0.01,\,0.1\}.$ (42)
For each candidate $\alpha$, models were trained on the training split and selected by the best validation metric. Final performance is reported on the held-out test set using the selected $\alpha$.
The optimal exploration parameter $\alpha$ for the LinUCB coincided with the optimal $\alpha$ for PSI-LinUCB -LinUCB algorithm.
A.1.3 Rank grids for PSI-LinUCB -LinUCB and $m$ for CBSCBD
Model capacity for both the PSI-LinUCB -LinUCB rank and the CBSCBD parameter $m$ was explored over a following grid :
$\text{Ranks}=\bigl\{\,b\cdot 2^{\,i}\;\big|\;i\in\{1,2,3,4,5\},\;b\in\{2,3\}\,\bigr\}.$ (43)
Selection followed the same validation protocol as for $\alpha$.
Table 3: Performance comparison for the increasing number of arms with rank $r$
Dataset Context Items Algorithm Hit Rate Train Time (s) Train Mem (MB) Eval Time (s) Eval Mem (MB)
Magazine d=41 (1681) r=64 1000 PSI-LinUCB -LinUCB 0.5957 52.33 1593.5 13.95 1726.0
LinUCB 0.5957 1367.94 43173.8 40.09 43519.1
LinUCB Classic 0.5957 446.70 21761.9 39.85 21934.4
2000 PSI-LinUCB 0.5533 57.99 1748.9 28.81 1901.7
LinUCB 0.5533 1743.39 84730.6 90.61 86585.7
LinUCB Classic 0.5533 485.78 42597.7 89.63 43525.8
3000 PSI-LinUCB 0.5482 59.77 1794.4 44.21 1971.3
LinUCB 0.5482 1848.22 124381.9 139.33 129707.1
LinUCB Classic 0.5482 503.49 62436.4 139.69 65113.4
Health d=37 (1369) r=128 1000 PSI-LinUCB 0.0296 271.36 3189.5 13.56 3718.5
LinUCB 0.0296 1719.74 28496.6 20.48 29239.3
LinUCB Classic 0.0296 844.69 14514.1 21.05 14885.6
10000 PSI-LinUCB 0.0189 512.89 6683.4 311.16 7726.5
LinUCB failed - - - -
LinUCB Classic 0.0189 1627.97 133827.8 713.53 143940.4
60000 PSI-LinUCB 0.0167 639.88 9662.3 2576.1 11146.8
LinUCB failed - - - -
LinUCB Classic failed - - - -
Beauty d=51 (2601) r=128 1000 PSI-LinUCB 0.0401 447.50 6290.4 16.47 7313.7
LinUCB 0.0401 10575.42 100768.9 57.98 103970.1
LinUCB Classic 0.0401 4636.82 50695.9 56.71 52297.3
70000 PSI-LinUCB 0.0142 1265.85 24840.5 5507.90 28724.0
LinUCB failed - - - -
LinUCB Classic failed - - - -
110000 PSI-LinUCB 0.0122 1328.77 27205.7 9189.7 31562.6
LinUCB failed - - - -
LinUCB Classic failed - - - -
Musical Instruments d=70 (4900) r=256 1000 PSI-LinUCB 0.2574 3858.76 32758.2 811.49 34460.6
LinUCB failed - - - -
LinUCB Classic 0.2574 75289.6 183968.5 4089.5 185250.7
80000 PSI-LinUCB 0.0606 13652.76 161977.2 131747.78 186052.1
LinUCB failed - - - -
LinUCB Classic failed - - - -
200000 PSI-LinUCB 0.0458 14535.0 192015.1 189541.5 201018.1
LinUCB failed - - - -
LinUCB Classic failed - - - -
Figure 6: Algorithm scaling with number of arms on Health & Personal Care dataset.
Figure 7: Algorithm scaling with number of arms on Magazine Subscriptions dataset.
Appendix B Proofs
B.1 Proof for Section˜4.2
Theorem 3.
Let $X_{t+1}\in\mathbb{R}^{d\times B}$ be a rank-$B$ update matrix with $B\ll d$. Then the updated matrix $A_{t}+X_{t+1}X_{t+1}^{\top}$ can be symmetrically factored as
$A_{t+1}=L_{t+1}L_{t+1}^{\top},$ (44)
where
$L_{t+1}=L_{t}(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}),$ (45)
and the factors are obtained as follows:
  1. 1.
    Compute $\bar{X}_{t+1}=L_{t}^{-1}X_{t+1}\in\mathbb{R}^{d\times B}$ and perform the QR decomposition
    $\bar{X}_{t+1}=Q_{t+1}R_{t+1},$ (46)
    where $Q_{t+1}\in\mathbb{R}^{d\times B}$ is an orthogonal matrix ($Q_{t+1}^{\top}Q_{t+1}=I_{B}$), and $R_{t+1}\in\mathbb{R}^{B\times B}$ is an upper triangular matrix.
  2. 2.
    Form the small matrix
    $T=I_{B}+R_{t+1}R_{t+1}^{\top}\in\mathbb{R}^{B\times B}.$ (47)
  3. 3.
    Compute the Cholesky decomposition of the small matrix
    $T=MM^{\top},$ (48)
    where $M\in\mathbb{R}^{B\times B}$ is a lower triangular matrix.
  4. 4.
    Set
    $Y_{t+1}=M-I_{B}\in\mathbb{R}^{B\times B}.$ (49)
This result follows from Remark 3.4 in Ambikasaran , see equations (3.5)–(3.6).
B.2 Proof of Theorem˜2
Proof.
Writing the Cholesky decomposition $A_{t}=L_{t}L_{t}^{\top}$ and factoring out $L_{t}$ gives
$\begin{split}A_{t+1}&=A_{t}+X_{t+1}X_{t+1}^{\top}\\[4.0pt] &=L_{t}L_{t}^{\top}+X_{t+1}X_{t+1}^{\top}\\[4.0pt] &=L_{t}\bigl(I+L_{t}^{-1}X_{t+1}X_{t+1}^{\top}L_{t}^{-T}\bigr)L_{t}^{\top}\\[4.0pt] &=L_{t}\bigl(I+\bar{X}_{t+1}\bar{X}_{t+1}^{\top}\bigr)L_{t}^{\top},\end{split}$
where $\bar{X}_{t+1}=L_{t}^{-1}X_{t+1}$. Applying Theorem 3, we obtain
$I+\bar{X}_{t+1}\bar{X}_{t+1}^{\top}=\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{\top},$ (50)
which yields
$A_{t+1}=L_{t}\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{\top}L_{t}^{\top}=L_{t+1}L_{t+1}^{\top},$ (51)
where
$L_{t+1}=L_{t}\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr).$ (52)
The inverse of $L_{t+1}$ is
$L_{t+1}^{-1}=\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{-1}L_{t}^{-1}.$ (53)
Applying the Sherman-Morrison-Woodbury formula gives
$\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{-1}=I-Q_{t+1}\bigl(Y_{t+1}^{-1}+I\bigr)^{-1}Q_{t+1}^{\top}.$ (54)
We proceed by induction to maintain a low-rank representation of $L_{t}^{-1}$. Assume that
$L_{t}^{-1}=(I-U_{t}V_{t}^{\top})L_{0}^{-1}.$ (55)
Then
$L_{t+1}^{-1}$ $=\bigl(I-Q_{t+1}(Y_{t+1}^{-1}+I)^{-1}Q_{t+1}^{\top}\bigr)L_{t}^{-1}$ (56)
$=\bigl(I-Q_{t+1}(Y_{t+1}^{-1}+I)^{-1}Q_{t+1}^{\top}\bigr)(I-U_{t}V_{t}^{\top})L_{0}^{-1}$ (57)
$=\bigl(I-U_{t}V_{t}^{\top}-Q_{t+1}(Y_{t+1}^{-1}+I)^{-1}Q_{t+1}^{\top}(I-U_{t}V_{t}^{\top})\bigr)L_{0}^{-1}$ (58)
$=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1},$ (59)
where
$U_{t+1}=\begin{bmatrix}U_{t}&Q_{t+1}(Y_{t+1}^{-1}+I)^{-1}\end{bmatrix}$ (60)
$V_{t+1}=\begin{bmatrix}V_{t}&(I-V_{t}U_{t}^{\top})Q_{t+1}\end{bmatrix}$ (61)
This completes the inductive step. ∎