### v2

# Scalable LinUCB: Low-Rank Design Matrix Updates for Recommenders with Large Action Spaces

Abstract
In this paper, we introduce PSI-LinUCB, a scalable variant of LinUCB that enables efficient training, inference, and memory usage by representing the inverse regularized design matrix as a sum of a diagonal matrix and low-rank correction. We derive numerically stable rank-1 and batched updates that maintain the inverse without explicitly forming the matrix. To control memory growth, we employ a projector-splitting integrator for dynamical low-rank approximation, yielding an average per-step update cost and memory usage of $\mathcal{O}(dr)$ for approximation rank $r$. The inference complexity of the proposed algorithm is $\mathcal{O}(dr)$ per action evaluation. Experiments on recommender system datasets demonstrate the effectiveness of our algorithm.
Keywords: LinUCB algorithm, Low-rank approximation, Projector-Splitting integrator
1 Introduction and Problem Setting
Contextual bandits are essential in modern decision making, as they enable online adaptation to dynamic environments and explicitly balance exploration and exploitation auer2002finite. These methods are well studied theoretically auer2002finite, abbasi2011improved, agrawal2013thompson and are widely used in practice, in particular, in the real-world recommender systems li2010contextual, yi2023online. Within this class of methods, LinUCB li2010contextual, abbasi2011improved is one of the most commonly used algorithms. It models the expected reward of an action as a linear function of a $d$-dimensional context vector and follows the Upper Confidence Bound (UCB) principle, implementing optimism in the face of uncertainty auer2002finite. LinUCB admits practical implementations for online recommendation settings li2010contextual and serves as a natural baseline when online updates are required.
For the most part of this paper, we focus on the disjoint LinUCB parametrization li2010contextual, das2024linear, yet the techniques that we develop further naturally generalize to hybrid and shared parametrizations used in this algorithm. Formally, at each time step $t\in\{1,\ldots,T\}$, agent observes a set of arms (actions) $\mathcal{A}_{t}$, where each arm $a\in\mathcal{A}$ is associated with a context vector ${x}_{t,a}\in\mathbb{R}^{d}$. LinUCB assumes a linear reward model,
$\mathbb{E}[r_{t,a}|{x}_{t,a}]={x}_{t,a}^{\top}{\theta}_{a}^{*},$ (2)
where ${\theta}_{a}^{*}\in\mathbb{R}^{d}$ is the unknown true parameter vector for arm $a$. For each arm $a\in\mathcal{A}$, the algorithm maintains a ridge regression estimator with a design matrix
$A_{t,a}=\lambda I+\sum_{s=1}^{t-1}{x}_{s,a}{x}_{s,a}^{\top},$ (3)
reward vector
$b_{t,a}=\sum_{s=1}^{t-1}r_{s,a}{x}_{s,a},$ (4)
and parameter estimate
$\hat{{\theta}}_{a}=A_{t,a}^{-1}b_{t,a}.$ (5)
At time $t$, the algorithm selects the arm according to
$a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\hat{{\theta}}_{a}^{\top}{x}_{t,a}+\alpha\sqrt{{x}_{t,a}^{\top}A_{t,a}^{-1}{x}_{t,a}}\right),$ (6)
where $\alpha\in\mathbb{R}$ controls the exploration–exploitation trade-off. In online learning problems $\alpha>0$, but sometimes learning problems with fixed dataset require setting negative values of $\alpha$ (the so-called anti-exploration), see e.g. rezaeifar2022offline, xu2024provably.
LinUCB can also be extended to batch-update settings, where model parameters are updated after multiple interactions (see Appendix˜A in Appendix). At the same time, LinUCB algorithm has several scalability challenges:
  • Matrix inversion: The complexity of computing $A_{t,a}^{-1}$ is $O(d^{3})$, which makes the algorithm computationally expensive if context dimension $d$ is large.
  • Large action space: The need to store and update a separate $d\times d$ matrix for each action leads to increased time and memory requirements as the number of actions grows.
Existing work addresses the scalability limitations of LinUCB in several ways. The simplest acceleration relies on exact rank-one updates of the inverse design matrix using the Sherman–Morrison formula Angioli2025Efficient, which avoids full matrix inversion but still requires maintaining dense $d\times d$ matrices and does not scale to large feature dimensions and large action space. Another line of work focuses on approximating the design matrix $\sum_{s=1}^{t}x_{s,a}x_{s,a}^{\top}$ with a low-rank representation using matrix sketching techniques kuzborskij2019efficient, chen2020efficient, wen2024matrix. While these methods significantly reduce memory and computation costs, they typically process observations sequentially and do not naturally support vectorized or batch updates. An alternative approach applies random feature projections to reduce the dimensionality of the context vectors before learning yu2017cbrap. Such methods enable faster updates but often require a relatively large projected dimension to preserve recommendation quality, making them less efficient than sketching-based low-rank approximations. Finally, diagonal or block-diagonal approximations of the design matrix have been proposed to reduce computational and memory costs yi2023online. We provide a detailed overview of the existing literature in the Appendix, Appendix˜B.
Our contrubutions. In this paper, we propose a scalable variant of LinUCB designed for large-scale contexts settings. Our approach is based on a suitable representation of the inverse regularized design matrix. Specifically, we approximate $A_{t,a}^{-1}$ as the sum of a diagonal matrix and a low-rank correction. This representation avoids explicit matrix inversion, and significantly reduces memory usage by storing only the diagonal terms and low-rank factors. Our primary contributions are as follows:
  • We introduce PSI-LinUCB, a scalable LinUCB variant that maintains a Cholesky-style representation of the inverse regularized design matrix as a sum of diagonal term and dynamically updated low-rank component. Using a projector-splitting method lubich2014projector, our algorithm avoids explicit matrix inversion. Our method naturally supports vectorized and batch updates over multiple context vectors, which is crucial for efficient deployment in modern recommender systems. PSI-LinUCB achieves an average per-interaction update cost of $\mathcal{O}(dr)$ under the proper choice of batch size, with $r$ being the approximation rank. The inference complexity is $\mathcal{O}(dr)$ per action.
  • We empirically show that PSI-LinUCB is robust with respect to the choice of the approximation rank: increasing $r$ leads to a smooth and nearly linear improvement in recommendation quality (measured by hit rate), making the method easy to tune in practice.
  • We evaluate PSI-LinUCB on several large-scale datasets from recommender systems domain, where standard LinUCB fails to run due to scalability constraints. We ensure reduction in memory consumption, and computational time compared to the exact Sherman–Morrison implementation of LinUCB Angioli2025Efficient, while matching the hit rate of the vanilla LinUCB algorithm. We also show that PSI-LinUCB outperforms sketching-based baselines in terms of computational time, while achieving similar or better quality, both on large-scale datasets and in online synthetic environments.
2 Scalable LinUCB with Low-Rank Updates
The main computational cost of LinUCB lies in updating the parameter estimate $\hat{\theta}_{a}$ for each arm, which requires access to the inverse matrix $A_{t,a}^{-1}$. Existing approaches mitigate this issue either by applying rank-1 Sherman–Morrison updates Angioli2025Efficient or by approximating the covariance matrix
$\sum_{s=1}^{t-1}{x}_{s,a}{x}_{s,a}^{\top}\in\mathbb{R}^{d\times d}\;,$ (7)
using low-rank sketching techniques such as Frequent Directions kuzborskij2019efficient and CBSCFD chen2020efficient. In contrast to these works, we employ an alternative approach, which relies on the dynamic representation of the inverse matrix $A_{t,a}^{-1}$. Towards this aim, we use the Cholesky-style symmetric factorization $A_{t,a}=L_{t,a}L_{t,a}^{\top}$ and note that in this case
$A_{t,a}^{-1}=L_{t,a}^{-\top}L_{t,a}^{-1}\;.$ (8)
The motivation for working with the inverse design matrix is that in LinUCB the design matrix $A_{t,a}$ is never used directly: both the parameter estimate
$\hat{\theta}_{t,a}=A_{t,a}^{-1}b_{t,a}$ (9)
and the exploration bonus
$\sqrt{x_{t,a}^{\top}A_{t,a}^{-1}x_{t,a}}$ (10)
involve only matrix–vector products with $A_{t,a}^{-1}$. Therefore, maintaining or approximating $A_{t,a}$ itself is algorithmically unnecessary.
The Cholesky-style factorization (8) makes this observation particularly convenient. Indeed, it allows both quantities above to be computed via matrix-vector operations involving $L_{t,a}^{-1}$ only. This representation enables us to work directly with the inverse operator without explicitly forming or storing $A_{t,a}^{-1}$.
The representation (8) is a key formula for our further analysis. Importantly, we do not explicitly compute the inverse factors $L_{t,a}^{-1}$. Instead, they are updated dynamically from their previous values. Since the update rules are identical across arms, we omit the index $a$ in the remainder of this section and write $A_{t}$ and $L_{t}$ (respectively, $L_{t}^{-1}$) instead of $A_{t,a}$ and $L_{t,a}$.
In the next parts of this section, we present our methodology in different setups. First, in Section˜2.1 we consider the case of rank-1 updates and write the dynamics of updates of the inverse root matrix $L_{t}^{-1}$. Then in Section˜2.2 we generalize our expressions for the batch update case. Both representations rely on a recursive definition in the form of matrix decomposition $U_{t}V_{t}^{\top}$. In Section˜3, we show that this decomposition can be dynamically updated in a low-rank format using the Projector-Splitting Integrator (PSI) method of Lubich and Oseledets lubich2014projector. Such updating mechanism prevents an uncontrollable growth of factor matrices $U_{t}$ and $V_{t}$ over time and enables “on-the-fly” adaptation to the stream of contextual data. This approach has been previously considered in the context of modifying the PureSVD model in recommender systems Olaleke2021Dynamic.
2.1 Rank-1 updates
We first consider the rank-1 update setting, where context vectors arrive sequentially and model parameters are updated after each interaction. In this case, the design matrix update at time $t+1$ takes the form
$A_{t+1}=A_{t}+x_{t+1}x_{t+1}^{\top}.$ (11)
While inverse updates can be obtained via the Sherman–Morrison formula Angioli2025Efficient, we instead derive an alternative representation that enables efficient low-rank updates within our proposed framework.
We begin by expressing the rank-1 update in a Cholesky-style factorized form. Using $A_{t}=L_{t}L_{t}^{\top}$, the update at time $t+1$ can be written as
$A_{t+1}=L_{t}L_{t}^{\top}+x_{t+1}x_{t+1}^{\top}=L_{t}(I+L_{t}^{-1}x_{t+1}x_{t+1}^{\top}L_{t}^{-\top})L_{t}^{\top}.$ (12)
Hence, we can rewrite the design matrix as
$A_{t+1}=L_{t+1}L_{t+1}^{\top}\;,$ (13)
where we set
$L_{t+1}=L_{t}(I+\alpha_{t+1}\tilde{x}_{t+1}\tilde{x}_{t+1}^{\top})\;,\;\tilde{x}_{t+1}=L_{t}^{-1}x_{t+1}\;,$ (14)
and parameter $\alpha_{t+1}\in\mathbb{R}$ such that
$1+\alpha_{t+1}\,\|\tilde{x}_{t+1}\|^{2}\;=\;\sqrt{\,1+\|\tilde{x}_{t+1}\|^{2}\,}\;.$ (15)
The correctness of the specified recursive definition of $L_{t+1}$ in (14) can be verified by direct substitution. The following theorem for the rank-1 updates setting will be useful for illustrating the essence of our approach.
Theorem 1.
Let $\varepsilon>0$, $L_{0}^{-1}=\varepsilon^{-1/2}I\in\mathbb{R}^{d\times d}$, and $U_{0}$, $V_{0}$ be empty matrices. Given a sequence of context vectors $\{x_{t}\}_{t\in\mathbb{N}}$ set
$\beta_{t+1}\;=\;\frac{\alpha_{t+1}}{1+\alpha_{t+1}\,\|\tilde{x}_{t+1}\|^{2}}.$ (16)
Then the inverse root $L_{t+1}^{-1}$ can be expressed as
$L^{-1}_{t+1}\;=\;\bigl(I-U_{t+1}\,V_{t+1}^{\top}\bigr)\,L_{0}^{-1},$ (17)
where the matrices $U_{t+1}\in\mathbb{R}^{d\times(t+1)}$ and $V_{t+1}\in\mathbb{R}^{d\times(t+1)}$ are recursively updated with column-wise concatenation
$U_{t+1}=\begin{bmatrix}U_{t}&\beta_{t+1}\,\tilde{x}_{t+1}\end{bmatrix},V_{t+1}=\begin{bmatrix}V_{t}&(I-V_{t}\,U_{t}^{\top})\,\tilde{x}_{t+1}\end{bmatrix}$. (18)
Proof.
The proof is provided in the Appendix, Section˜D.1
2.2 Batch updates
Now we provide a generalization of Theorem˜1 to the case of batch updates. In this setting parameter is updated after some number of interactions is accumulated in the system. Formally, we write $X_{t}\in\mathbb{R}^{d\times B}$ for the batch of $B\in\mathbb{N}$ concatenated contexts collected during the $t$-th interaction round with the arm $a\in\mathcal{A}$. Note that, generally speaking, $B$ depends on $a$ and $t$, but we prefer to write $B$ instead of $B_{t,a}$ for notation simplicity.
The update rule for the regularized design matrix can then be expressed as
$A_{t+1}=A_{t}+X_{t+1}X_{t+1}^{\top}=L_{t+1}L_{t+1}^{\top},$ (19)
where the Cholesky-like factor $L_{t+1}$ is obtained recursively:
$L_{t+1}=L_{t}(I+Q_{t+1}(M_{t+1}-I)Q_{t+1}^{\top}),$ (20)
with $Q_{t+1}\in\mathbb{R}^{d\times B}$ being a matrix with orthonormal columns and $M_{t+1}\in\mathbb{R}^{B\times B}$ obtained using the fast symmetric factorization approach Ambikasaran described in Appendix˜D (see Theorem˜3).
Theorem 2.
Let $\varepsilon>0$, $L_{0}=\varepsilon^{-1/2}I\in\mathbb{R}^{d\times d}$ and $U_{0},V_{0}$ be empty matrices, and $U_{t},V_{t}\in\mathbb{R}^{d\times d_{t}}$. Then the factor $L_{t+1}$ defined in (20) can be updated by formula
$L_{t+1}^{-1}=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1}.$ (21)
Matrices $U_{t+1},V_{t+1}\in\mathbb{R}^{d\times(d_{t}+B)}$ are updated with column-wise concatenation
[…]
Proof of Theorem˜2 is provided in Section˜D.3. The shape $d_{t}$ of the matrices $U_{t}$ and $V_{t}$ can be inferred from the representations (21) - (2).
2.3 Low-rank correction for $A_{t}^{-1}$
Note that the matrices $U_{t}$ and $V_{t}$ in (2) expand in the number of columns with time $t$, leading to excessive memory demands and computational overhead. We now motivate why the correction term for $\varepsilon^{-1}I-A_{t}^{-1}$ can be well approximated by a low-rank matrix. Consider the empirical covariance matrix $\sum_{s=1}^{t}x_{s}x_{s}^{\top},$ and let its SVD be given by
$\sum_{s=1}^{t}x_{s}x_{s}^{\top}=E_{x}\Sigma_{x}E_{x}^{\top}.$ (22)
Then the regularized design matrix and its inverse admit the representations
$A_{t}=E_{x}(\varepsilon I+\Sigma_{x})E_{x}^{\top},\qquad A_{t}^{-1}=E_{x}(\varepsilon I+\Sigma_{x})^{-1}E_{x}^{\top}.$ (23)
Then the correction term $R_{t}:=\varepsilon^{-1}I-A_{t}^{-1}$ writes as
$R_{t}=E_{x}\Bigl(\varepsilon^{-1}I-(\varepsilon I+\Sigma_{x})^{-1}\Bigr)E_{x}^{\top}.$ (24)
The diagonal entries of the matrix inside the parentheses are given by
$\varepsilon^{-1}-\frac{1}{\varepsilon+\sigma_{i}}=\frac{\sigma_{i}}{\varepsilon(\varepsilon+\sigma_{i})}\leq\frac{\sigma_{i}}{\varepsilon^{2}}.$ (25)
Therefore, eigencomponents corresponding to small eigenvalues $\sigma_{i}$ contribute negligibly to $R_{t}$. Moreover, approximating the empirical covariance $\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ with rank $r$ before inversion $A_{t}$ or approximating only the correction $R_{t}$ with rank $r$ in decomposition for $A_{t}^{-1}$ leads to the same approximation of $A_{t}^{-1}$. This observation supports modeling $R_{t}$ using a low-rank representation when the empirical covariance matrix $\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ has low effective rank, a property commonly employed in the literature on sketching methods kuzborskij2019efficient, chen2020efficient. Below we describe how to maintain a low-rank approximation of $R_{t}$ by controlling the ranks of $U_{t}$ and $V_{t}$ using the Projector-Splitting Integrator lubich2014projector. The motivation for this particular approximation method is discussed in the next section.
3 Projector-Splitting Integrator
We now aim to improve LinUCB using the representation in (21)–(2). In this form, the matrices $U_{t}$ and $V_{t}$ grow with time. At the same time, the key element of the representation (21) is the time-dependent matrix $D_{t}=U_{t}V_{t}^{\top}\in\mathbb{R}^{d\times d}$, whose evolution defines the dynamics of the representation. A natural approach is to approximate $D_{t}$ via a rank-$r$ truncated SVD with,
$D_{t}=\bar{U}_{t}\Sigma_{t}\bar{V}_{t}^{\top}\;.$ (26)
At the same time, the orthogonal factors $(\bar{U}_{t},\bar{V}_{t})$ and $(\bar{U}_{t+1},\bar{V}_{t+1})$, corresponding to the matrices $D_{t}$ and $D_{t+1}$, are not guaranteed to be close, which might yield additional computational instability. Instead, we propose to rely on the projector-splitting integrator (PSI) approach of lubich2014projector, which constructs a dynamical low-rank approximation by solving
$\|\dot{D}_{t}-\dot{\tilde{D}}_{t}\|\to\min_{\tilde{D}_{t}:\,\operatorname{rank}{\tilde{D}}_{t}=r}\,$, (27)
where $\dot{D}_{t}=\frac{\mathrm{d}}{\mathrm{d}t}D_{t}$. The PSI method enables efficient iterative updates without explicitly forming $\tilde{D}_{t}$. The approximation is maintained in factorized form,
$\tilde{D}_{t}=\tilde{U}_{t}S_{t}\tilde{V}_{t}^{\top}\;,$ (28)
where the factors $\tilde{U}_{t},\tilde{V}_{t}\in\mathbb{R}^{d\times r}$ have orthonormal columns, and $S_{t}\in\mathbb{R}^{r\times r}$ is invertible. Then we obtain factors $(\tilde{U}_{t+1},S_{t+1},\tilde{V}_{t+1}^{\top})$, such that
$\tilde{D}_{t+1}=\tilde{U}_{t+1}S_{t+1}\tilde{V}_{t+1}^{\top}\;.$ (29)
We provide implementation details in Algorithm˜3.
We integrate this procedure into LinUCB with batch updates as follows. During training, the factors $U_{t}$ and $V_{t}$ are incrementally expanded using (2) or (44) until their number of columns exceed a predefined threshold $r$. At this time $t_{0}$, we compute a rank-$r$ SVD of $U_{t_{0}}V_{t_{0}}^{\top}$ to initialize the PSI factors. Subsequent updates apply Algorithm˜3 to maintain a fixed-rank approximation. The update increments $\Delta D_{t+1}$ are obtained directly from the LinUCB update rules. For rank-one updates,
$\Delta D_{t+1}=\beta_{t+1}\tilde{x}_{t+1}\tilde{x}_{t+1}^{\top}(I-U_{t}V_{t}^{\top});$ (30)
and from (2) for the batch update case:
$\Delta D_{t+1}=Q_{t+1}(M_{t+1}-I)M_{t+1}^{-1}Q_{t+1}^{\top}(I-U_{t}V_{t}^{\top}).$ (31)
The complete integration of PSI into LinUCB is summarized in Algorithm˜1. We highlight that we do not need to form the matrices $L_{t}^{-1}$ given in (21) explicitly.
Complexity analysis. We analyze the computational complexity of the proposed PSI-LinUCB algorithm during training. The complexity of the PSI update
$U_{t+1},V_{t+1}=\textit{PSI}(U_{t},V_{t},\;\Delta D_{t+1})$ (32)
does not depend on the number of interactions $B=B_{t,a}$ (that is, interactions with arm $a$ inside batch number $t$) and scales as $\mathcal{O}(dr^{2})$. This factor comes from the QR decomposition applied to $\mathbb{R}^{d\times r}$ matrices. When the rank exceeds the threshold, an SVD is computed once per arm (line 11 in Algorithm˜2) with complexity $\mathcal{O}(d(r+B)^{2}+(r+B)^{3})$, obtained via QR decompositions of factors of size at most $\mathbb{R}^{d\times(r+B)}$. Additionally, line 5 in Algorithm˜2 incurs a cost of $\mathcal{O}(dB^{2}+B^{3})$ due to QR and Cholesky decompositions. Thus, the overall complexity of handling blocks in the Update_arm algorithm (Algorithm˜2) is
$\mathcal{O}(d(r^{2}+B^{2})+r^{3}+B^{3})\;,$ (33)
as soon as this arm has accumulated at least $r$ interactions.
Inference stage. To compute the bonus term in (6) we use the decomposition defined earlier:
$\sqrt{x_{t,a}^{\top}A_{t}^{-1}x_{t,a}}$ $=\sqrt{x_{t,a}^{\top}(L_{t}L_{t}^{\top})^{-1}x_{t,a}}$ (34)
$=\|(I-U_{t}V_{t}^{\top})L_{0}^{-1}x_{t,a}\|$
Then at time $t$, the arm is selected by formula:
$\begin{split}a_{t}=\arg\max_{a\in\mathcal{A}_{t}}\left(\theta_{a}^{\top}x_{t,a}+\alpha\|(I-U_{t}V_{t}^{\top})L_{0}^{-1}x_{t,a}\|\right)\;.\end{split}$
The computational complexity of this operation is $\mathcal{O}(dr)$ per arm.
Table 1: Computational complexity comparison across algorithms
Algorithm Time cost per round Space
LinUCB $O(d^{2})$ $O(d^{2})$
CBRAP $O(dm+m^{3})$ $O(dm)$
CBSCFD $O(dm)$ $O(dm)$
DBSLinUCB $O(dl_{B_{t}})$ $O(dl_{B_{t}})$
PSI-LinUCB $O(dr)$ $O(dr)$
Remark 1.
The primary computational cost of our algorithm arises during the training phase, which requires on average $\mathcal{O}(d(r^{2}/B+B)+r^{3}/B+B^{2})$ operations per interaction. By choosing $B\propto r$ and assuming $r\ll d$, this reduces to average complexity of $\mathcal{O}(dr)$ operations per iteration. To our knowledge, the $\mathcal{O}(dr)$ average complexity is the best computational cost achieved by LinUCB-based algorithms. This is comparable with the most sample-efficient implementations of existing algorithms chen2020efficient, wen2024matrix, summarized in Table 1.
Theoretical properties. Assume that the matrix $U_{t}V_{t}^{\top}$ from the exact LinUCB update representation Theorem˜2 has rank not exceeding $r$ for any $t\in\{0,\ldots,T\}$. Then theoretical guarantees for PSI integrator applied with the same rank $r$ lubich2014projector[Theorem 4.1] ensures exact integration, that is, Algorithm˜3 maintains the exact $L_{t}^{-1}$ and and hence the exact inverse $A_{t}^{-1}=L_{t}^{-\top}L_{t}^{-1}$. Consequently, the estimator $\hat{\theta}_{t}=A_{t}^{-1}b_{t}$ and the confidence bonus in (34) exactly matches those of the standard LinUCB algorithm (6). In this setting, PSI-LinUCB admits the standard regret scaling of LinUCB under linear bandit assumptions of order $\tilde{O}(d\sqrt{T})$ with high probability after $T$ iterates, see abbasi2011improved.
Investigating the setting when the exact LinUCB algorithm matrix $U_{t}V_{t}^{\top}$ has larger rank is an important direction for future work. Existing regret analysis of LinUCB and its sketching variants kuzborskij2019efficient rely on the monotonicity properties of the estimates of feature covariance matrix. In our setting, we directly approximate the inverse regularized design matrix, and such monotonicity does not hold, making standard techniques unavailable.
Algorithm 1 PSI-LinUCB (LinUCB training with PSI)
0: train_data = $batch_{0},\dots,batch_{n-1}]$, batch_size, U_0,V_0U\_0,\,V\_0
[…]
2:for t in {0,…,n-1} do
3:  $X_{t,a}=[]$, $R_{t,a}=[]$
4:  for all $(u,a,r)$ in $batch_{t}$ do
5:   $X_{t,a}.append\,[x_{t,a}]$, $R_{t,a}.append\,[r]$
6:  end for
7:  for each arm $a$ in $batch_{t}$ do
8:   $U_{t+1,a},V_{t+1,a},\theta_{t+1,a}\leftarrow$ Update_arm($a$)
9:  end for
10:end for
11:return $\theta_{t,a}$, $U_{t,a},V_{t,a}$ for each arm $a\in\mathcal{A}$.
Algorithm 2 Update_arm($a$)
[…]
1: Set $b_{t}=b_{t,a}$, $U_{t}=U_{t,a}$, $V_{t}=V_{t,a}$, $X_{t}=X_{t,a}$, $R_{t}=R_{t,a}$, {Omitting index $a$ for simplicity}
[…]
6:if $U_{t}.\text{shape}[1]<r$ then
7:  $U_{t+1}\;\leftarrow\;[\,U_{t},\;Q_{t+1}C_{t+1}]$
8:  $V_{t+1}\;\leftarrow\;[\,V_{t},\,(I-V_{t}U_{t}^{\top})\;Q_{t+1}\,]$
9:else
10:  if first time $U_{t}.\text{shape}[1]\geq r$ then
11:   $\tilde{U}_{t+1}S_{t+1}\tilde{V}_{t+1}^{\top}=SVD(U_{t}V_{t}^{T})$
12:   $U_{t+1}\;\leftarrow\;\tilde{U}_{t+1}S_{t+1}$
13:   $V_{t+1}\;\leftarrow\;\tilde{V}_{t+1}$
14:  else
15:   $\Delta D_{t+1}=\,Q_{t+1}\,C_{t+1}\,Q_{t+1}^{\top}\,(I-U_{t}V_{t}^{\top}),$ {We do not form $\Delta D_{t+1}$ explicitly}
16:   $U_{t+1},V_{t+1}=\textit{PSI}(U_{t},V_{t},\;\Delta D_{t+1})$
17:  end if
18:end if
[…]
21:return $U_{t+1},V_{t+1}$, $\theta_{t+1}$
Algorithm 3 PSI(Projector-Splitting Integrator)
[…]
Algorithm 4 Calculate_C_and_Q
[…]
4 Experimental Setup
We evaluate the proposed PSI-LinUCB against five baselines covering exact updates, sketching and random-projection approaches:
  • LinUCB li2010contextual: the standard LinUCB algorithm with batched updates;
  • LinUCB Classic  Angioli2025Efficient: LinUCB with exact rank-1 inverse updates via the Sherman–Morrison formula;
  • CBSCFD  chen2020efficient: a sketching-based method that approximates the feature covariance (design) matrix.
  • CBRAP yu2017cbrap: method based on random projections applied to feature covariance (design) matrix.
  • DBSLinUCB wen2024matrix: an adaptive sketching method that dynamically adjusts the sketch size over time.
We provide more details about existing methods in the Appendix, Appendix˜B.
Table 2: Information about the datasets used for validation
Dataset #Users #Items #Interactions Density
Magazine Subscriptions 60,100 3,400 71,500 0.35%
Health & Personal Care 461,700 60,300 494,100 0.02%
All Beauty 632,000 112,600 701,500 0.01%
MovieLens 1M 6,040 3,706 1,000,209 4.18%
4.1 Training Protocol
In order to run our algorithm on the datasets, we convert them into the bandit style environment following the pipeline below:
  • Warm-up phase: Initial training on 80% of historical data
  • Online training phase: Sequential learning on the remaining 20% of data, divided into equal-sized temporal intervals simulating monthly updates. The model is incrementally trained on each month’s data.
  • Evaluation: Model evaluated after each monthly training update to assess recommendation quality over time.
In the sections below, we consider the datasets described in Table˜2. The LinUCB algorithm is implemented with batch processing to ensure scalability with respect to high-dimensional contextual features. User-item features are extracted via SVD decomposition of the user-item interaction matrix, with rank $r^{\prime}$ selected through spectral analysis during warm-up. To obtain high-dimensional contexts, the context vectors are constructed as outer products of user and item embeddings. All hyperparameters are tuned via cross-validation for each dataset and model, see Section˜C.1 in Appendix.
5 Scalability results
Increasing context size. To ensure the scalability of PSI-LinUCB, we vary the dimensions of context vectors $x_{t,a}$. We infer user-item features from SVD with varying rank $r^{\prime}$ on the subset of the Amazon Health dataset, and measure the computational time and memory usage. As shown in Figure˜1, LinUCB and LinUCB Classic fail to scale beyond moderate context sizes due to memory and time requirements for storing and inverting the full matrix $A_{t,a}$. In contrast, PSI-LinUCB operates efficiently even for large $d$ by maintaining only low-rank factors with $\mathcal{O}(dr)$ complexity. As shown in Table˜3, PSI-LinUCB achieves identical quality to LinUCB where both are feasible, while continuing to operate at larger context sizes. Warm-up phase results are provided in Appendix C, see Figure˜9.
Figure 1: Performance comparison across different context sizes on Amazon Health dataset.
Table 3: Hit Rate comparison across context dimensions $d$
$d$ LinUCB LinUCB Woodbury PSI-LinUCB
900 0.026 0.026 0.026
2500 0.034 0.034 0.034
4900 - - 0.044
16900 - - 0.050
Increasing number of arms. We fix the optimal context dimensions and PSI rank for each dataset (Health, Beauty, and Magazine Subscriptions), then gradually increase the number of arms from a small subset to the full action space, selecting items with the highest observation frequency at each stage. As shown in Figure˜2, exact implementations of LinUCB become grossly memory-inefficient as number of arms increases, while PSI-LinUCB remains computationally tractable. Additional results on time and memory consumption during the warm-up phase on the Beauty dataset (Figure 10(c)), as well as further scalability experiments are reported in the Appendix,  Appendix˜C.
Figure 2: Algorithm scaling with number of arms on Beauty dataset.
LinUCB quality. We analyze how the rank parameter affects recommendation quality by varying the PSI rank while keeping context dimensions and number of arms fixed. At each configuration, we measure hit rate of our algorithm, popular and random baselines relative to the one of LinUCB. Figure˜3 shows that PSI-LinUCB achieves recommendation quality comparable to classical LinUCB starting from relatively low ranks, without requiring a full-rank matrix. Additional results are provided in Appendix, see Figure˜8.
(a) Amazon Health
(b) Amazon All Beauty
Figure 3: Quality comparison by months: PSI-LinUCB vs LinUCB on Amazon datasets.
6 Comparison with scalable baselines
6.1 Online setting
Following the experimental setup of kuzborskij2019efficient and chen2020efficient, we evaluate our method on an online classification problem turned into contextual bandit setting. To maintain compatibility with the mentioned works, in this subsection we switch to the version of LinUCB with shared parametrization, see e.g. chen2020efficient with a single parameter vector and design matrix across all arms.
Given a dataset with $K$ classes, we mark of them as a target. In each round, the environment randomly draws one sample from each class, forming a context set. The learner selects one sample and receives a reward of $1$ if the selected sample belongs to the target cluster, and 0 otherwise. Note that in this binary reward setting, cumulative number of mistakes corresponds to the cumulative regret. We conduct experiments on two benchmark datasets: MNIST LeCun1998Gradient and CIFAR-10 Krizhevsky2009Learning. For fair comparison we tuned the optimal rank/sketch size for all algorithms via cross-validation by minimizing mistakes over $T$ rounds. The dataset statistics are summarized in the Appendix, Table˜6, the detailed description of the experimental setup and hyperparameter selection is provided in Section˜6.3. Since this experiment focuses on online learning, we use the rank-1 version of PSI-LinUCB (see Algorithm˜6 in Appendix), which is consistent with the update scheme employed by CBSCFD chen2020efficient, CBRAP yu2017cbrap, and DBSL wen2024matrix.
The results, presented in Figure 4, demonstrate that our PSI-LinUCB algorithm achieves competitive performance with CBSCFD on MNIST while being significantly faster and outperforms all other baseline methods in both quality and computational efficiency. Results on CIFAR dataset are presented in Appendix, Appendix˜C, Figure 11).
Figure 4: Online classification results on the MNIST dataset.
6.2 Approximation quality of the inverse matrix
Figure 5 shows the dependence of the relative approximation error of the inverse design matrix $A_{t}^{-1}$ on final iteration for different ranks (sketch sizes for CBSCFD). The approximation error of PSI-LinUCB and CBSCFD decreases as the rank increases. However, PSI-LinUCB consistently achieves lower error than CBSCFD for the same rank size. The corresponding experiments on CIFAR (see Figure 16) and real-world datasets (see Figures 16(a), 16(b)) are presented in Appendix,  Appendix˜C.
Figure 5: Approximation error of $A_{t}^{-1}$ on MNIST
6.3 Performance on real-world datasets
This experiment compares PSI-LinUCB with CBSCFD, CBRAP and DBSLinUCB across multiple datasets. In the original methods, a shared design matrix $A_{t}$ is used for all arms; we extend this by training a separate matrix $A_{t,a}$ for each arm $a$, as proposed in li2010contextual. Additionally, CBSCFD and CBRAP support only rank-1 updates, while PSI-LinUCB generalizes to batch updates. The experiment follows the setup described in Section˜4.1, for full details see Section˜C.1 in Appendix.
Table 4 presents quality metrics across three datasets (Amazon Magazine Subscriptions, Amazon All Beauty, Amazon Health), and Table 5 reports training time, per-user prediction time and memory usage. CBRAP achieves quality comparable to PSI-LinUCB and CBSCFD but requires a larger sketch size $m$, resulting in higher memory and time costs. This explains inherently large training and evaluation time of CBRAP on the Beauty dataset. DBSLinUCB introduces additional tuning complexity, requiring joint search over the initial block size $l_{0}$ (which grows exponentially) and the error parameter $\varepsilon$, with interdependent effects. For example, DBSLinUCB fails on the Health dataset with all checked hyperparameters, see Table˜4. CBSCFD shows comparable quality to PSI-LinUCB, yet we point out its main drawback as instability of the quality with respect to sketch size $m$.
Figure˜6 and Figure˜7 show the relationship between the parameter $m$ and hit rate for CBSCFD and PSI-LinUCB, respectively. The results demonstrate significant variation in the optimal $m$ values across different real-world datasets, with unpredictable behavior: even small changes in $m$ lead to different results. At the same, for PSI-LinUCB the dependence of quality on approximation rank $r$ is monotone as the rank increases, which is desirable for tuning this (arguably, most important) hyperparameter in practice. We provide additional experiments on other datasets in the Appendix, Appendix˜C (see Figure 14).
Table 4: Quality metrics on different datasets
Dataset Algorithm Hit@10 NDCG@10 MRR@10 Cov.
Magazine PSI-LinUCB 0.531 0.430 0.398 0.038
CBSCFD 0.526 0.449 0.425 0.031
CBRAP 0.528 0.420 0.386 0.027
DBSLinUCB 0.523 0.438 0.411 0.028
Health PSI-LinUCB 0.022 0.017 0.016 0.0016
CBSCFD 0.020 0.015 0.014 0.0012
CBRAP 0.020 0.017 0.016 0.0012
DBSLinUCB 0.001 0.001 0.001 0.0006
Beauty PSI-LinUCB 0.013 0.011 0.010 0.0076
CBSCFD 0.012 0.009 0.009 0.0011
CBRAP 0.012 0.011 0.010 0.0011
DBSLinUCB 0.013 0.011 0.010 0.0047
Table 5: Computational efficiency on different datasets
Dataset Algorithm Train (s) Eval (s) Train (GB) Eval (GB)
Magazine PSI-LinUCB 10.3 0.145 1.4 1.6
CBSCFD 22.9 0.163 1.5 1.5
CBRAP 134.9 9.224 1.5 1.5
DBSLinUCB 24.6 0.102 1.5 1.8
Health PSI-LinUCB 25.9 1.24 6.9 8.8
CBSCFD 61.8 2.28 8.4 8.9
CBRAP 563.5 96.37 10.9 16.6
DBSLinUCB 518.8 1.5 7.3 9.1
Beauty PSI-LinUCB 83.6 3.5 21.7 26.8
CBSCFD 106.5 6.01 22.4 27.2
CBRAP 1566.0 350.0 41.7 67.0
DBSLinUCB 290.8 3.6 19.1 24.8
6.4 Batch size and rank trade-off
The batch size $B$ controls the frequency of PSI updates. Larger batches reduce training time but may slightly degrade quality. Figures 17(a) and 17(b) in Appendix show that training time decreases substantially with larger batches while hit rate remains stable. This enables flexible tuning: larger batches for latency-sensitive deployments, smaller batches for quality-critical scenarios.
(a) Amazon Health
(b) Amazon All Beauty
Figure 6: Average Hit Rate for different $m$ for CBSCFD.
(a) Amazon Health
(b) Amazon All Beauty
Figure 7: Average Hit Rate for different $r$ for PSI-LinUCB.
7 Conclusion
We presented PSI-LinUCB, a scalable variant of LinUCB for large-scale contextual bandits. The method maintains a compact representation of the inverse regularized design matrix using a diagonal term and a low-rank correction, which allows efficient computation. Moreover, our method has an average complexity of updates of order $\mathcal{O}(dr)$ per candidate. A key direction for future work is to develop theoretical guarantees under direct approximations of the inverse regularized design matrix. To our knowledge, there are no regret guarantees for linear bandit algorithms in this setting. At the same time, as we show numerically, such algorithms might be preferable as compared to the classical sketching algorithms in particular applications.
Acknowledgements
This research was supported in part through computational resources of HPC facilities at HSE University kostenetskiy2021hpc.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A LinUCB with Batch Updates
Algorithm 5 LinUCB
0: train_data = $batch_{0},\dots,batch_{n-1}]$, regularization parameter $\lambda$
1: for each arm $a\in\mathcal{A}$ set $A_{0,a}=\lambda I$
2:for t in {0,…,n-1} do
3:  $X_{t,a}=[]$, $R_{t,a}=[]$
4:  for all $(u,a,r)$ in $batch_{t}$ do
5:   $X_{t,a}.append\,[x_{t,a}]$
6:   $R_{t,a}.append\,[r]$
7:  end for
8:  for each arm $a$ in $batch_{t}$ do
9:   $A_{t+1,a}\leftarrow A_{t,a}+X_{t,a}^{\top}X_{t,a}$
10:   $b_{t+1,a}\leftarrow b_{t,a}+X_{t,a}^{\top}R_{t,a}$
11:   ${\theta}_{t+1,a}\leftarrow A_{t+1,a}^{-1}b_{t+1,a}$
12:  end for
13:end for
14:for each arm $a\in\mathcal{A}$ do
15:  Observe context $x_{t,a}$
16:  Compute $\text{UCB}_{t,a}={\theta}_{t,a}^{\top}{x}_{t,a}+\alpha\sqrt{{x}_{t,a}^{\top}{A}_{t,a}^{-1}{x}_{t,a}}$
17:end for
18:return $\arg\max_{a\in\mathcal{A}}\text{UCB}_{t,a}$
Appendix B Related Work and Baseline Selection
B.1 Scalable variants of LinUCB
Although LinUCB li2010contextual is a widely used algorithm in recommendation systems, the time and memory requirements increases with the dimension of the context $d$ and the number of items, since it is necessary to store and invert the matrix $A_{a}\in\mathbb{R}^{d\times d}$ for each item $a$. To address scalability constraints, several extensions and variants of the LinUCB algorithm have been proposed in the literature.
A widely used approach to accelerate LinUCB is to apply rank-1 updates of the ridge-regularized design matrix $A_{a}$ via the Sherman–Morrison identity Angioli2025Efficient, Ciucanu2022Implementing, Wang2022Dynamic, Yan2025CoCoB, Ozbay2024Comparative, zenati2022efficient. These updates reduce the per-round update complexity from $\mathcal{O}(d^{3})$ to $\mathcal{O}(d^{2})$ while maintaining exact parameter estimates, but still require storing a full $d\times d$ matrix for each arm.
Another line of research aims to improve the efficiency of LinUCB through matrix sketching techniques. Early work kuzborskij2019efficient demonstrated that LinUCB can be efficiently implemented using the Frequent Directions (FD) sketching method, reducing the per-round update time from $\mathcal{O}(d^{2})$ to $\mathcal{O}(md)$, where $m$ is the sketch size and $d$ is the feature dimension. However, applying FD to contextual bandits has certain drawbacks. In particular, the FD sketching method violate the positive definite monotonicity design matrices $A_{t,a}$, which may affect stability and theoretical guarantees. To address this limitation, a Spectral Compensation Frequent Directions (SCFD) method and its adaptation (CBSCFD) for high-dimensional contextual bandit were proposed chen2020efficient, which preserves positive definiteness while maintaining the same $\mathcal{O}(md)$ computational and memory complexity. More recently, adaptive sketching techniques such as Dyadic Block Sketching wen2024matrix have been introduced to dynamically adjust the sketch size, ensuring a per-round update complexity of $\mathcal{O}(dl)$, where $l$ is the current sketch size. This adaptive strategy prevents excessive spectral loss and avoids linear regret when the spectrum of the design matrix decays slowly.
Another perspective on improving the efficiency of linear contextual bandits is through dimensionality reduction via random projection. For example, the Contextual Bandits via Random Projection (CBRAP) algorithm yu2017cbrap, address the challenges of high-dimensional contexts by mapping the original $d$-dimensional features to a lower $m$-dimensional subspace. This reduces the per-round update complexity from $\mathcal{O}(d^{2})$ to $\mathcal{O}(md+m^{3})$.
One more approach to tackle the limitations of the LinUCB algorithm is to approximate each design matrix $A_{a}$ along its diagonal as proposed in Diag-LinUCB yi2023online. This allows scalable online updates running as $\mathcal{O}(d)$ per round in terms of both time and memory. As the authors demonstrate, in the specific contexts $x_{ua}$ such updates are sufficient to prevent the loss of model quality during training.
B.2 Baselines used in our experiments
To evaluate PSI-LinUCB, we select baselines that cover both exact implementations and implementations using covariance matrix approximation, which reduce either the effective rank of the matrix or the feature dimension.
As exact implementations, we report results for Batch LinUCB [li2010contextual], which matches batched training protocol provided in Algorithm˜5 and LinUCB Classic with Sherman–Morrison rank-one inverse updates [Angioli2025Efficient]. These baselines provide a useful quality reference and illustrate the memory and computational cost of exact updates in high dimensions.
To represent design-matrix compression via sketching, we use CBSCFD [chen2020efficient], which combines FD with an additional correction and maintains a low-rank sketch of the design matrix. CBSCFD is widely used as a robust sketching baseline for high-dimensional linear contextual bandits. We additionally include Dyadic Block Sketching (DBSLinUCB) [wen2024matrix] as a recent adaptive sketching approach that dynamically adjusts sketch size.
To cover feature-space compression, we include CBRAP [yu2017cbrap], which applies random projections before performing LinUCB-style updates. This baseline is conceptually different from sketching the design matrix and provides a typical alternative for dimensionality reduction.
Some methods are closely related but are not included as primary baselines. We do not include FD [kuzborskij2019efficient] in the main comparison because CBSCFD [chen2020efficient] combines FD with an additional correction and provides a stronger and more robust representative within the same sketching family. We also omit Diag-LinUCB [yi2023online] because it suggests to learn context representations (using non-linear transformations) followed by diagonal approximation and thus the comparison with original LinUCB is uninformative.
Appendix C Experiments
C.1 Hyperparameter Selection
Context dimensionality $d$
The optimal values of $r^{\prime}$ were selected using the scree plot, which displays singular values in descending order. The resulting context sizes are:
  • MovieLens 1M: $r^{\prime}=13$, $d=169$;
  • Magazine Subscriptions: $r^{\prime}=41$, $d=1681$;
  • Health & Personal Care: $r^{\prime}=37$, $d=1369$;
  • All Beauty: $r^{\prime}=51$, $d=2601$.
Regularization $\alpha$
We used grid search to find the optimal exploration parameter $\alpha$. First, a broad grid $\alpha\in\{-10,-8,\dots,10\}$ was evaluated, followed by a refined search over $\alpha\in\{-1,-0.5,-0.3,-0.2,-0.1,-0.05,-0.03,-0.01,0.1\}$. The optimal $\alpha$ for LinUCB coincided with that for PSI-LinUCB.
Rank and sketch size
PSI-LinUCB rank, CBSCFD/CBRAP parameter $m$, DBSLinUCB block size $l_{0}$ (which is not fixed but grows exponentially during learning) and the error parameter $\varepsilon$, was explored over the grid $\{b\cdot 2^{i}\mid i\in\{1,\dots,8\},b\in\{2,3\}\}$.
Feature extraction
Given an interaction matrix $A\in\mathbb{R}^{m\times n}$, we apply truncated SVD: $A\approx U\Sigma V^{\top}$. Item features are obtained from $V^{\top}\in\mathbb{R}^{n\times r}$, and user features from $A\cdot V^{\top}\in\mathbb{R}^{m\times r}$. Features are extracted from warm-up data, cold users lacking valid context representations are excluded from evaluation. Note that cold-start users is not inherent to bandits, as soon as there are available contexts, the algorithm can be run, the restriction arises from our feature construction SVD-based pipeline and does not alter the fundamental bandit framework. Then we construct a context ${x}_{t,a}=\text{vec}(x_{u}{x}_{a}^{\top})\in\mathbb{R}^{d_{u}\times d_{a}}$, where $x_{u}$ and $x_{a}$ are user and item feature vectors. In our experiments, we set $d_{u}=d_{a}=r^{\prime}$, yielding $d=(r^{\prime})^{2}$.
Data processing
Training time measures initial warm-up training; evaluation time is the average prediction duration per user, averaged across test months. The interaction dataset is grouped by arms, with each batch corresponding to a fixed number of interactions for a single arm. This scheme is applied uniformly across all algorithms for fair comparison. The batch-to-rank ratio (see Remark 1) in PSI-LinUCB corresponds to the compression frequency factor of 2 in CBSCFD; however, batch size does not affect final model state or quality for CBSCFD and CBRAP, as these methods perform incremental per-sample updates.
(a) MovieLens 1M
(b) Magazine Subscriptions
Figure 8: Average Hit Rate for different $r$ for PSI-LinUCB.
Figure 9: Performance comparison across different context sizes on Amazon Health dataset.
(a) Health & Personal Care
(b) Magazine Subscriptions
(c) Beauty
Figure 10: Algorithm scaling with number of arms on different datasets.
Figure 11: Online classification results on CIFAR-10.
Figure 12: Approximation error of $A^{-1}$ on CIFAR-10.
Table 6: Summary of datasets for online classification
Dataset #Samples #Features #Classes
MNIST 60000 784 10
CIFAR-10 50000 3072 10
(a) MovieLens 1M
(b) Amazon All Beauty
(c) Amazon Health
(d) Magazine Subscriptions
Figure 13: Average Hit Rate for different $m$ values for CBRAP.
(a) MovieLens 1M
(b) Magazine Subscriptions
Figure 14: Average Hit Rate for different $m$ for CBSCFD.
(a) MovieLens 1M
(b) Magazine Subscriptions
Figure 15: Average Hit Rate for different rank values for PSI-LinUCB.
(a) Amazon Health
(b) Magazine Subscriptions
Figure 16: Relative approximation error of $A^{-1}$ on real-world datasets.
C.2 Setting of online classification
The hyperparameters $\beta$ and $\lambda$ are selected via grid search over $\{10^{-4},10^{-3},\ldots,1\}$ and $\{2\times 10^{-4},2\times 10^{-3},\ldots,2\times 10^{4}\}$, respectively, following the experimental protocol of chen2020efficient. For our PSI-LinUCB algorithm, the rank parameter $r$ and for sketching-based methods (CBSCFD, CBRAP) the optimal sketch size $m$ are selected via cross-validation over a grid ranging from $10$ to $100$ with step $10$ and $\{200,300,400,500\}$ by minimizing the number of the average online mistakes on a validation run.
All experiments are repeated $20$ times and we report the average cumulative mistakes.
This dual applicability represents a practical advantage of our approach: the same algorithmic framework can be deployed both in the per-arm setting commonly used in real-world recommendation systems and in the shared model setting typical of experimental benchmarks. This allows for detailed theoretical analysis in controlled settings while maintaining direct applicability to production environments.
In addition to the baselines from chen2020efficient, we include a comparison with the Dyadic Block Sketching Linear (DBSL) algorithm wen2024matrix, which represents a more recent adaptive sketching approach.
Algorithm 6 PSI-LinUCB Rank-1 Update
0: context $x_{t}$, reward $r_{t}$, rank $r$
[…]
6:if $U_{t}.\text{shape}[1]<2r$ then
7:  $U_{t+1}\leftarrow[U_{t},\;\beta_{t+1}\bar{x}_{t+1}]$
8:  $V_{t+1}\leftarrow[V_{t},\;(I-V_{t}U_{t}^{\top})\bar{x}_{t+1}]$
9:end if
10:if $U_{t}.\text{shape}[1]=2r$ then
11:  if first time $U_{t}.\text{shape}[1]=2r$ then
12:   $\tilde{U}_{t+1}S_{t+1}\tilde{V}_{t+1}^{\top}=SVD(U_{t}V_{t}^{\top})$
13:  end if
14:  $\Delta D_{t+1}=\beta_{t+1}\tilde{x}_{t+1}\tilde{x}_{t+1}^{\top}(I-U_{t}V_{t}^{\top})$ {We do not form $\Delta D_{t+1}$ explicitly}
15:  $\tilde{U}_{t+1},S_{t+1},\tilde{V}_{t+1}=\textit{PSI}(\tilde{U}_{t},S_{t},\tilde{V}_{t},\Delta D)$
16:  $U_{t+1}\;\leftarrow\;\tilde{U}_{t+1}S_{t+1}$
17:  $V_{t+1}\;\leftarrow\;\tilde{V}_{t+1}$
18:end if
[…]
20:return $U_{t+1},V_{t+1},\theta_{t+1}$
(a) Amazon Health
(b) Amazon All Beauty
Figure 17: Batch/rank dependence on Amazon datasets.
Appendix D Proofs
D.1 Proof of Theorem˜1
Using (14), we obtain
$L_{t+1}^{-1}=(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})^{-1}L_{t}^{-1}\;.$ (35)
After applying the Sherman-Morrison-Woodbury formula
$(I+\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})^{-1}=I-\frac{\alpha_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}}{1+\alpha_{t+1}\bar{x}_{t+1}^{\top}\bar{x}_{t+1}}=I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}\;,$ (36)
the recursive structure of the inverse update reads:
$L_{t}^{-1}=(I-\beta_{t}\bar{x}_{t}\bar{x}_{t}^{\top})L_{t-1}^{-1}\;.$ (37)
Unraveling this recursion from $t$ back to the initial condition gives
$L_{t}^{-1}=(I-\beta_{t}\bar{x}_{t}\bar{x}_{t}^{\top})L_{t-1}^{-1}=\prod_{i=1}^{t}(I-\beta_{i}\bar{x}_{i}\bar{x}_{i}^{\top})L_{0}^{-1}.$ (38)
We now prove by induction that $L^{-1}_{t}=(I-U_{t}\,V_{t}^{\top})L^{-1}_{0}$. Indeed, for $t=1$, we have
$L_{1}^{-1}=(I-\beta_{1}\bar{x}_{1}\bar{x}_{1}^{\top})L_{0}^{-1}=(I-U_{1}V_{1}^{\top})L_{0}^{-1}\;,$ (39)
where $U_{1}=\beta_{1}\bar{x}_{1}$ and $V_{1}=\bar{x}_{1}$. Assuming that $L^{-1}_{t}=(I-U_{t}\,V_{t}^{\top})L^{-1}_{0}$,
$L_{t+1}^{-1}$ $=(I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})L_{t}^{-1}$ (40)
$=(I-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top})(I-U_{t}\,V_{t}^{\top})L^{-1}_{0}$ (41)
$=(I-U_{t}V_{t}^{\top}-\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}+\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}U_{t}V_{t}^{\top})L_{0}^{-1}$ (42)
$=(I-(U_{t}V_{t}^{\top}+\beta_{t+1}\bar{x}_{t+1}\bar{x}_{t+1}^{\top}(I-U_{t}V_{t}^{\top})))L_{0}^{-1}\;.$ (43)
This can be written as $L_{t+1}^{-1}=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1}$, where
$U_{t+1}$ $=\begin{bmatrix}U_{t}&\beta_{t+1}\,\bar{x}_{t+1}\end{bmatrix}$, (44)
$V_{t+1}$ $=\begin{bmatrix}V_{t}&(I-V_{t}U_{t}^{\top})\bar{x}_{t+1}\end{bmatrix}$. (45)
This completes the proof.
D.2 Proof for Section˜2.2
Theorem 3.
Let $X_{t+1}\in\mathbb{R}^{d\times B}$ be a rank-$B$ update matrix with $B\ll d$. Then the updated matrix $A_{t}+X_{t+1}X_{t+1}^{\top}$ can be symmetrically factored as
$A_{t+1}=L_{t+1}L_{t+1}^{\top},$ (46)
where
$L_{t+1}=L_{t}(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}),$ (47)
and the factors are obtained as follows:
  1. 1.
    Compute $\bar{X}_{t+1}=L_{t}^{-1}X_{t+1}\in\mathbb{R}^{d\times B}$ and perform the QR decomposition
    $\bar{X}_{t+1}=Q_{t+1}R_{t+1},$ (48)
    where $Q_{t+1}\in\mathbb{R}^{d\times B}$ is an orthogonal matrix ($Q_{t+1}^{\top}Q_{t+1}=I_{B}$), and $R_{t+1}\in\mathbb{R}^{B\times B}$ is an upper triangular matrix.
  2. 2.
    Form the small matrix
    $T_{t+1}=I_{B}+R_{t+1}R_{t+1}^{\top}\in\mathbb{R}^{B\times B}.$ (49)
  3. 3.
    Compute the Cholesky decomposition of the small matrix
    $T_{t+1}=M_{t+1}M_{t+1}^{\top},$ (50)
    where $M\in\mathbb{R}^{B\times B}$ is a lower triangular matrix.
  4. 4.
    Set
    $Y_{t+1}=M_{t+1}-I_{B}\in\mathbb{R}^{B\times B}.$ (51)
This result follows from Remark 3.4 in Ambikasaran, see equations (3.5)–(3.6).
D.3 Proof of Theorem˜2
Proof.
Writing the Cholesky decomposition $A_{t}=L_{t}L_{t}^{\top}$ and factoring out $L_{t}$ gives
$\begin{split}A_{t+1}&=A_{t}+X_{t+1}X_{t+1}^{\top}\\[4.0pt] &=L_{t}L_{t}^{\top}+X_{t+1}X_{t+1}^{\top}\\[4.0pt] &=L_{t}\bigl(I+L_{t}^{-1}X_{t+1}X_{t+1}^{\top}L_{t}^{-T}\bigr)L_{t}^{\top}\\[4.0pt] &=L_{t}\bigl(I+\bar{X}_{t+1}\bar{X}_{t+1}^{\top}\bigr)L_{t}^{\top},\end{split}$
where $\bar{X}_{t+1}=L_{t}^{-1}X_{t+1}$. Applying Theorem 3, we obtain
$I+\bar{X}_{t+1}\bar{X}_{t+1}^{\top}=\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{\top},$ (52)
which yields
$A_{t+1}=L_{t}\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{\top}L_{t}^{\top}=L_{t+1}L_{t+1}^{\top},$ (53)
where
$L_{t+1}=L_{t}\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr).$ (54)
The inverse of $L_{t+1}$ is
$L_{t+1}^{-1}=\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{-1}L_{t}^{-1}.$ (55)
Applying equation (51) and the Sherman-Morrison-Woodbury formula gives
$\bigl(I+Q_{t+1}Y_{t+1}Q_{t+1}^{\top}\bigr)^{-1}$ $=I-Q_{t+1}\bigl((M_{t+1}-I_{B})^{-1}+I\bigr)^{-1}Q_{t+1}^{\top}$ (56)
$=I-Q_{t+1}(M_{t+1}-I)\bigl(M_{t+1}-I+I\bigr)^{-1}Q_{t+1}^{\top}$ (57)
$=I-Q_{t+1}(M_{t+1}-I)M_{t+1}^{-1}Q_{t+1}^{\top}$ (58)
We proceed by induction to maintain a low-rank representation of $L_{t}^{-1}$. Assume that
$L_{t}^{-1}=(I-U_{t}V_{t}^{\top})L_{0}^{-1}.$ (59)
Then
$L_{t+1}^{-1}$ $=\bigl(I-Q_{t+1}(M_{t+1}-I)M_{t+1}^{-1}Q_{t+1}^{\top}\bigr)L_{t}^{-1}$ (60)
$=\bigl(I-Q_{t+1}(M_{t+1}-I)M_{t+1}^{-1}Q_{t+1}^{\top}\bigr)(I-U_{t}V_{t}^{\top})L_{0}^{-1}$ (61)
$=\bigl(I-U_{t}V_{t}^{\top}-Q_{t+1}(M_{t+1}-I)M_{t+1}^{-1}Q_{t+1}^{\top}(I-U_{t}V_{t}^{\top})\bigr)L_{0}^{-1}$ (62)
$=(I-U_{t+1}V_{t+1}^{\top})L_{0}^{-1},$ (63)
where
$U_{t+1}=\begin{bmatrix}U_{t}&Q_{t+1}(M_{t+1}-I)M_{t+1}^{-1}\end{bmatrix}$ (64)
$V_{t+1}=\begin{bmatrix}V_{t}&(I-V_{t}U_{t}^{\top})Q_{t+1}\end{bmatrix}$ (65)
This completes the inductive step. ∎